### v1

# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities in natural language processing tasks. However, their application in the biomedical domain presents unique challenges, particularly regarding factual accuracy and up-to-date knowledge integration. Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these challenges by combining the generative capabilities of LLMs with external knowledge retrieval. This comprehensive survey examines the application of RAG in the biomedical domain, focusing on its technological components, available datasets, and clinical applications. We present a systematic analysis of retrieval methods, ranking strategies, and generation models, while also exploring the challenges and future directions in this rapidly evolving field. Our work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems and identifies key areas for future research and development.
Keywords
Biomedical RAG, Large Language Model, Retrieval-Augmented Generation, Information Retrieval, Natural Language Processing, Biomedical NLP
Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 53, 81. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 219.
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 128. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 227. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 58. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 53, 81, 227.
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 230. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 131, 184.
Concurrent to our survey, several related surveys focused on biomedical-related LLMs. For example, Liu et al. 135 review medical LLMs and He et al. 81 discuss LLMs for healthcare. Similarly, Xiao 227 conducted a comprehensive overview of LLMs in medicine. Other discussed RAG systems in general domains, such as Fan et al. 53 and Gao et al. 65 focused on the architectures, training strategies, and application of RAG in general domains.
Our work differs from existing surveys in concentrating on the application of RAG in the biomedical domain, systematically reviewing its components, datasets, and challenges. Overall, this survey provides the following key contributions:
  1. 1.
    A systematic analysis of RAG technologies specifically adapted for biomedical applications, including specialized retrieval methods, ranking algorithms, and generation models.
  2. 2.
    A detailed examination of medical datasets, considering both knowledge source datasets and medical task datasets.
  3. 3.
    An exploration of current clinical applications and future directions, with particular attention to practical implementation challenges and emerging opportunities.
  4. 4.
    A structured framework for understanding the interplay between retrieval mechanisms, medical knowledge sources, and language models in biomedical RAG systems.
The remainder of this survey is organized as follows: Section 2 presents the background and the basic framework of biomedical RAG systems. Section 3 details the technical components, including retrieval methods, ranking algorithms, and generation models. Section 4 examines available datasets. Section 5 explores current clinical applications, and Section 6 discusses challenges and future directions.
Background
LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 247. Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 197. RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 65. This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 227.
In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources. The re-ranking method then sorts these references based on their relevance. Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response. As Fig. 1 shows, the basic framework of biomedical RAG consists of four key components 53: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
Figure 1: Basic biomedical RAG framework.
RAG has been applied to a variety of biomedical applications. In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 141, enhances specialized decision-making through expert-guided LLMs 119, and facilitates rare disease diagnosis 240. This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 64 and advances genomic medicine through improved variant annotation and interpretation 142. In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 177. Across these applications, RAG leverages diverse data sources—including health records 105, 83, scientific literature 37, clinical guidelines 21, and genomic databases 142—to generate evidence-based responses that significantly improve both clinical care and research outcomes. The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
Technology
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 99, 243 and dense retrieval methods 168, 76, LLM-based retrieval 128, 163, knowledge graph enhancement techniques 108, 237, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 199, 177, 215. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 97, 94, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 165, 18, 73 and open-source models 7, 204, 156, 44 alongside domain adaptation techniques 58, 19 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
Retriever
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 65. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 218, 203. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains190. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this section, we describe various retrieval methodologies. We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
Sparse Retrieval
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) 182 and TF-IDF (Term Frequency-Inverse Document Frequency) 106, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 42, 71 implementations to hybrid approaches 120, 13, 3, 115 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 109. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 234, 88, 92. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch 51, Whoosh 206, and Entrez API 56 also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 231, 120, 13, 3, 115, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
Table 1: Sparse retrieval methods in biomedical RAG systems
Method Name Retrieval Method Task Year
Clinfo.ai 141 Entrez API Medical QA 2023
MEDRAG toolkit 231 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS 88 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System 120 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA 13 BM25 and MedCPT Medical Question Answering 2024
CliniqIR 3 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery 115 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT 142 Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT 71 BM25 algorithm Medical QA 2025
Two-Layer RAG 42 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt 169 Entrez API for PubMed retrieval Scientific data analysis 2025
Dense Retrieval
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 246. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 190. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 183, and Sentence Transformer 207 have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding 228, Dense Passage Retriever 111 and GIST Large Embedding 193 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText 26, RoBERTa 137, and SimCSE 63, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
Table 2: Classification of dense retriever types in biomedical RAG systems
Method Retriever Task Year
Type 1: PLMs
AskFDALabel 19 Sentence Transformer 207 FDA drug labeling extraction 2024
CaLM 168 BGE-embedding 228 with Chroma DB Supporting caregivers FM 2024
RAG-HPO 67 FastEmbed 205 Rare genetic disorders automated deep phenotyping 2024
RALL 76 Dense Passage Retriever 111 Lay language generation 2024
RT 129 BERT embeddings 46 Few-shot medical NER 2024
RAG-GPT 233 BGE-embedding 228 Breast cancer nursing care QA 2024
GNQA 41 HNSW graphs 147 Medical QA 2024
Klang et al. 116 GIST Large Embedding 193 with FAISS ICD-10-CM coding 2024
RAGPR 248 SBERT embeddings 183 Personalized physician recommendations 2025
JGCLLM 61 GLuCoSE-base-ja vectors 60 Genetic counseling support 2025
Type 2: Commercial APIs
DRAGON-AI 209 OpenAI Text-Embedding 164 with Chroma DB Ontology generation 2024
RISE 217 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG 121 UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT 130 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat 59 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT 139 Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT 198 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER 214 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP 20 Amazon Titan Text Embeddings v2 14 Nutrition-related question answering 2025
InfectA-Chat 186 Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG 107 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
Type 3: Domain adaptation PLMs
PM-Search 99 BioBERT 123 Clinical literature retrieval 2022
LADER 98 PubMedBERT 75 Biomedical literature retrieval 2023
CLEAR 140 BioBERT Clinical information extraction 2025
KG-RAG 194 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule 6 MedCPT, BioBERT, BioGPT 145 with Pinecone DB Biomedical QA 2024
Type 4: LLM-based embedding
BiomedRAG 128 MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM 163 GPT4All 16 with Chroma DB Surgical decision support 2024
Myers et al. 160 LLM2Vec (for comparison) 22 Clinical Information retrieval 2024
Type 5: Domain adaptation retriever
Self-BioRAG 93 MedCPT 97 Medical QA 2024
WeiseEule 6 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR 3 MedCPT and BM25 Diagnostic decision support 2024
MEDRAG 6 BM25, SPECTER 37, Contriever 92and MedCPT Biomedical RAG Tool 2024
RAMIE 242 MedCPT, Contriever and BMRetriever 234 Biomedical IR about dietary supplements 2025
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings dominate, alternatives are emerging. The RAP system 20 utilizes Amazon’s Titan Text Embeddings v2 14 for nutrition-related question answering. Other commercial APIs such as Anthropic’s Voyage-2 17 and Google’s Vertex AI embeddings 74, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 123, 75. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 146. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 146. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 75, BioBERT 123 and BioGPT 145, which have been widely adopted in biomedical RAG systems. For example, PM-Search 99 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER 98 for biomedical literature retrieval and in KG-RAG 194 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule 6 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 97 to achieve enhanced performance in biomedical QA.
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 247. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 128, which utilizes MedLLaMA 13b 224, SurgeryLLM 163, which employs GPT4All 16, and the comparative study by Myers et al. 160, which evaluates LLM2Vec 22 for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 138, Landmark Embedding 143, which provide promising dense retrieval alternatives in biomedical RAG.
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 92, MedCPT 97, SPECTER 37, BMRetriever 234, and the MEDRAG toolkit 128. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 93, 3, 242. For example, Jeong et al. 93 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework 6 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 242. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems 80. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3. In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
Table 3: Common vector stores in biomedical RAG
Store Description
Chroma 36 AI-native open-source vector database with built-in functionality
Faiss 52 Library for efficient similarity search and clustering of dense vectors
Pinecone 172 Production-ready vector database for similarity search at scale
Weaviate 220 Graph-based vector search engine with semantic capabilities
HNSW 147 Hierarchical navigable small world graphs for approximate nearest neighbor search
Knowledge Graph Enhanced Retrieval
Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability171. Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source. As we can notice, the application areas vary significantly, including medical question answering 194, knowledge extraction 47, and diagnosis prediction 64.
Table 4: Recent biomedical RAG studies based on knowledge graphs
Method Task Knowledge Graph Used Year
KG-RAG 194 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE 108 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN 152 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle 237 Various medical text generation tasks UMLS 2024
DALK 125 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. 69 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 47 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO 229 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS 64 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT 151 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
As shown in Fig 2, the integration of knowledge graphs with LLMs in biomedical RAG applications can be analyzed through four key components:
  • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. NetMe 2.0 47 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 229 mines PubMed literature for synthetic biology knowledge graphs. Differently, KG-RAG 194 utilizes the pre-existing SPOKE biomedical knowledge graph.
  • Query to knowledge graph transformation encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. ESCARGOT 151 employs Cypher queries for precise retrieval, KRAGEN 152 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 64 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
  • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. DR.KNOWS 64 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 152 applies graph-of-thoughts for multi-hop relationships, and HEALIE 108 extracts paths incorporating both medical knowledge and socio-economic factors.
  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency. KG-RAG 194 implements token-optimized retrieval to reduce consumption, ESCARGOT 151 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 47 transforms graph data into natural language prompts via its Graph-RAG module.
Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
Multi-modal Retrieval
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 4. As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 215, prescription label interpretation 208, lung cancer staging 211, digital pathology analysis 162, and automated radiology report generation 178, 199, 225, 226, 238.
Figure 3: Architectural workflow of multimodal biomedical RAG system
Table 5: Multimodal Biomedical RAG Systems
Model Task Modalities Year
Ranjit et al. 178 Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG 199 Radiology report generation X-Ray and text 2024
MMed-RAG 225 Medical VQA & Report generation Medical images and text 2024
MEAG 215 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al.  211 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG 162 Digital pathology QA Pathology images and text 2024
Raminedi et al. 177 Radiology report generation X-ray images and text 2024
RULE 226 Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. 208 Prescription label identification Image (labels) and text 2025
Hu et al. 87 Pathology report generation Whole slide images and text 2025
STREAM 238 Chest X-ray report generation Chest X-ray and text 2025
Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 4. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 32 for images, specialized encoders for time-series data 32); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.178, FactMM-RAG 199 and STREAM 238 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG 225 and RULE 226 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG 215 system uses specialized encoders for eye movement data alongside text, Raminedi et al. 162 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. 208 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
Reranker
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 190. Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 3 and MEDRAG 6 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 209 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 6 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 94 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 5 utilizes a specialized cross-encoder model known as BGE M3-Embedding 31 to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 97 trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 41 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 39 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai 141 combines LLM classification with BM25 to re-rank retrieved articles.
Generation
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 122. This section examines various generation models employed in recent biomedical RAG studies.
Current generation models employed in biomedical RAG systems can be categorized into three distinct classes based on their development approach and specialization: general-domain open-source models, commercial proprietary models, and biomedical-specialized models. These categories present different trade-offs in terms of accessibility, customization potential, and domain-specific performance. Table LABEL:tab:llm_taxonomy showcases representative models from each category and highlights recent studies that utilize these models. It is important to note that our analysis of biomedical-specialized LLMs focuses specifically on models trained directly on biomedical or healthcare corpora, deliberately excluding applications developed through prompt engineering or Chain-of-Thought (CoT) methodologies.
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 247. Among open-source LLMs, LLaMa 2 210 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 7, Mistral 7B 155, and T5 176 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 156 and Gemma 3 72. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 217, 84, 119, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies 166, 242, 11 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
Table 6: Taxonomic Classification of LLMs for Embedding Applications
Model Parameters Architecture Studies
Open-Source LLMs
T5 176 0.06-11B Encoder-Decoder VAIV 115, CLEAR 140, DR.KNOWS 64
ChatGLM3 70 6B Decoder-only COPD 244, ChatZOC 144
LLaMA 2 210 7-70B Decoder-only Guo et al. 76, KREIMEYER et al. 117, CaLM 168, MMRAG 241, Yu et al. 239, RAMIE 242, EyeGPT 33, KG-RAG 194, JMLR 219, BiomedRAG 128, Self-BioRAG 93, CaLM 168, Du et al. 49, SurgeryLLM 163, Kreimeyer et al. 117
LLaMA 3 7 8-70B Decoder-only MMRAG 241, Fatharanihttps et al. 54, RAMIE 242, Woo et al. 222, RAG-HPO 67, i-MedRAG 232, Hewitt et al. 84, PodGPT 94, SurgeryLLM 163
Phi-3 Mini 154 3.8B Decoder-only Fatharanihttps et al. 54
Mistral 7B 155 7B Decoder-only Boulos et al. 110, RAMIE 242, Woo et al. 222, RAGPR 248, LITURAt 169, Kreimeyer et al. 117
QWen 175 32B Decoder-only NEKO 229, Thetbanthad 208
Baichuan 216 7-13B Decoder-only ChatZOC 144
Falcon 12 7-180B Decoder-only AskFDALabel 19, CaLM 168
Zephyr 213 7B Decoder-only Moser et al. 159
AceGPT 89 7-13B Decoder-only InfectA-Chat 186
Gemma 62 1-27B Decoder-only PodGPT 94
Deepseek R1 44 671B Decoder-only Feng et al. 55
Commercial LLMs
ChatGPT-3.5/4/4o 165 Proprietary Decoder-only Woo et al. 222, RAGPR 248, ChatZOC 144, KG-RAG 194, Clinfo.ai 141, Yu et al. 239, ChatENT 139, CaLM 168, Rau et al. 180, Endo-chat 59, BiomedRAG 128, Lu 142, Tan 15, GNQA 40, Puts 173, DRAGON-AI 209, Choi 35, RISE 217, Du et al. 49,i-MedRAG 232 , Hewitt et al. 84, RECTIFIER 214, Kainer et al. 107, Gong et al. 71, VAIA 115, RareDxGPT 240, Selcuk et al. 186, FAVOR-GPT 130, Markey et al. 148,
Claude-3.5 18 Proprietary Decoder-only Woo et al. 222, RISE 217, Hewitt et al. 84
Gemini 73 Proprietary Multimodal Upadhyaya et al. 215, Tozuka et al. 211, MEREDITH 119
Biomedical Specialized LLMs
MedAlpaca 79 7-13B LLaMA-based RAMIE 242
PMC-LLaMA 223 7B LLaMA-based RAMIE 242, BiomedRAG 128, Ozaki et al. 166
BioMistral 118 7B Mistral-based RAMIE 242, pRAGe 28
MEDITRON 34 7-70B LLaMA-based Ozaki et al. 166, Alkaeed et al. 11
Meerkat 113 7B Mistral-based Sohn et al. 192
Domain Adaptation
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 189. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 92 and MedCPT 97) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 234) and domain-specific architectural modifications (exemplified by LADER 98 with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 97) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 128). Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria. Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 58, SeRTS 88), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 19), and continued pre-training (RALL 76, ChatENT 139) enables more accurate synthesis and explication of biomedical information. JLMR 219 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns essential for healthcare and life sciences applications. Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, ultimately advancing the state-of-the-art in biomedical information retrieval and synthesis.
Datasets
The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets. In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task datasets. This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
Knowledge Source Datasets
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 66, 53. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 90, 139, 93.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries 6, 234.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS 25, PMC, and Medical textbooks 95 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies 35, 49, 248 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 120, 150, 233, 20, 89, 187 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
Table 7: Comprehensive Medical Knowledge Sources
Source Type Description Studies
MIMIC-IV 103 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 159
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 141, MEDRAG 231, Self-BioRAG 93, JMLR 219, Ascle 237, RISE 217, MEREDITH 119, LITURAt 169, DR.KNOWS 64, Aftab 6, CliniqIR 3, PM-Search 99, Kreimeyer 117, Kim 115
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 93, Soong et al. 196, PodGPT 94
TripClick 181 Literature Biomedical literature search and retrieval dataset LADER 98
GNQA 40 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 40
UMLS 25 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 128, CLEAR 140, Guo 76, CliniqIR 3
ICD-10 30 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 173
HPO 102 Knowledge Base Human phenotype ontology Albayrak et al. 9, RAG-HPO 67, Kainer 107
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 234
AlzKB 151 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 151, KRAGEN 152
DrugBank 221 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 234, Kim 115
SPOKE 158 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 194
COD 144 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 144
HEALIE KG108 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 108
ADRD Caregiving 168 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 168
StatPearls 21 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 231, i-MedRAG 232
Medical Textbooks 95 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 231, Self-BioRAG 93, JMLR 219, i-MedRAG 232
Ophthalmology textbooks 33 Educational 14 specialized ophthalmology textbooks EyeGPT 33
CheXpert 91 Image-Report 224,316 chest radiographs with associated reports CLEAR 140
Medical Task Datasets
Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 81, 135. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups 61, 198, 217 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
Table 8: Comprehensive Biomedical Tasks Datasets
Dataset Year Task Studies
ChemProt 202 2010 Information Extraction BiomedRAG 128, ChemProt
DDI 185 2013 Information Extraction BiomedRAG 128, ChemProt
NCBI 48 2014 Entity Recognition GeneGPT 100, NetMe 47, RT 129
BioASQ 212 2015 Information Retrieval BMRETRIEVER 234, SeRTS 88
MIMIC-III 105 2016 Text Summarization DR.KNOWS 64, CliniqIR 3
BC5CDR 126 2016 Entity Recognition RT 129
DC3 50 2019 Diagnostic classification CliniqIR 3
MIMIC-CXR 104 2019 Text Summarization Ascle 237
MedQuAD 23 2019 QA Ascle 237
PubMedQA 96 2019 Multiple-choice BMRETRIEVER 234, Liévin et al. 133, PodGPT 94
MMLU 82 2021 Multiple-choice Self-BioRAG 93, JMLR 219, i-MedRAG 232, InfectA-Chat 186, PodGPT 94
MedQA 95 2021 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, i-MedRAG 232, PodGPT 94
HealthSearchQA 188 2022 QA
MedMCQA 167 2022 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, EyeGPT 33, PodGPT 94
ChatDoctor 132 2023 Dialogue EyeGPT 33
MedAlpaca 79 2023 Dialogue EyeGPT 33
GeneTuring 86 2023 Genomics QA GeneGPT 100
GIT 127 2023 Information Extraction BiomedRAG 128
MIRAGE 231 2024 Multiple-choice & QA MEDRAG 231
MedExpQA 13 2024 Multilingual Medical QA MedExpQA 13, PodGPT 94
PubMedRS-200 141 2024 Medical QA Clinifo.ai 141
MedicineQA 90 2024 Multi-round dialogue RagPULSE 90
CELLS 76 2024 Lay language generation Guo 76
Applications
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 136. In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
Clinical Decision Support Systems
Medical Question Answering
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 194. This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai 141 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL 76 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information. Self-BioRAG 93 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 6 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
Diagnostic and Treatment Decision Support
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 201. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs 119 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS 64 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR 3 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 163 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER 214 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
Rare Disease Identification and Management
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 149. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. 240 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO 67 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. 236 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. Some authors 114 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
Clinical Report Generation
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 191. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 78. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 10.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 230. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 179, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors 8 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines 24 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG 195 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors 200 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
Precision Medicine Applications
Genomic Medicine
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 29.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT 130 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. 142 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO 67 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 61 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 142 and phenotypic and genotypic analysis 67, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
Personalized Health Management
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 174. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE 108 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE 217 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System 120 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP 20 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
Healthcare Education
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 124. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT 139 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT 33 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. 187 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval. In nursing education, Zhao et al 245 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.
Clinical Research
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. 68.
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT 94 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt 169 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
Challenges and Future Direction
Trustworthiness
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 53, 45. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 66. Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 122. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora 81, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
Multilinguality
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 161, 101. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 157, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
Multimodality
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray 104, 170, 134), time series (e.g. ECG 239), sound recordings (e.g. lung sounds recorded 57), and video 77. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 249. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the Biomedical domain and improving its accuracy and trustworthiness in the future.
Restricted Computing Resources
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 53. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 43.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 27, 85. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts 235. Models such as QWQ 175 and Mistral Small 3.1 156 demonstrate competitive performance despite having significantly reduced parameter counts. Applying Portable LLMs 112 as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
Privacy
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 153. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 38 and the Health Insurance Portability and Accountability Act 2. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.
Conclusion
In conclusion, this comprehensive survey has provided a detailed overview of the application of RAG in the biomedical domain. It has examined the basic framework of biomedical RAG and discussed its retriever, reranker, and generation modules, as well as explored KG-enhanced biomedical RAG and multi-modal biomedical RAG. The survey has also explored clinical applications where biomedical RAG can be applied and reviewed recent studies in this field. Additionally, the paper has identified key challenges, including trustworthiness, multilingual capabilities, multimodal integration, limited computational resources, and privacy concerns. Overall, this work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems, ultimately contributing to the development of more accurate, accessible, and impactful healthcare AI systems that can meaningfully enhance patient care through evidence-based information access and support.

### v2

@@
 Introduction
-Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 53, 81.
+Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 1, 2.
 These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval.
-The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 219.
+The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 3.
 However, the application of LLMs in the biomedical domain faces two critical challenges.
-First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 128.
+First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 4.
 This is particularly concerning in medical applications, where accuracy is paramount for patient safety.
-Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 227.
+Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 5.
 This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
-Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 58.
-By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 53, 81, 227.
+Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 6.
+By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 1, 2, 5.
 The implementation of RAG in the biomedical domain presents unique challenges and opportunities.
-Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 230.
-Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 131, 184.
+Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 7.
+Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 8, 9.
 Concurrent to our survey, several related surveys focused on biomedical-related LLMs.
@@
 2 discuss LLMs for healthcare.
-Similarly, Xiao 227 conducted a comprehensive overview of LLMs in medicine.
+Similarly, Xiao 5 conducted a comprehensive overview of LLMs in medicine.
 Other discussed RAG systems in general domains, such as Fan et al.
@@
 Background
-LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 247.
-Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 197.
-RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 65.
-This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 227.
+LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 12.
+Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 13.
+RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 11.
+This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 5.
 In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources.
@@
 As Fig.
-1 shows, the basic framework of biomedical RAG consists of four key components 53: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
+1 shows, the basic framework of biomedical RAG consists of four key components 1: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
 Figure 1: Basic biomedical RAG framework.
 RAG has been applied to a variety of biomedical applications.
-In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 141, enhances specialized decision-making through expert-guided LLMs 119, and facilitates rare disease diagnosis 240.
-This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 64 and advances genomic medicine through improved variant annotation and interpretation 142.
-In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 177.
-Across these applications, RAG leverages diverse data sources—including health records 105, 83, scientific literature 37, clinical guidelines 21, and genomic databases 142—to generate evidence-based responses that significantly improve both clinical care and research outcomes.
+In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 14, enhances specialized decision-making through expert-guided LLMs 15, and facilitates rare disease diagnosis 16.
+This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 17 and advances genomic medicine through improved variant annotation and interpretation 18.
+In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 19.
+Across these applications, RAG leverages diverse data sources—including health records 20, 21, scientific literature 22, clinical guidelines 23, and genomic databases 18—to generate evidence-based responses that significantly improve both clinical care and research outcomes.
 The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
@@
 In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information.
-Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 99, 243 and dense retrieval methods 168, 76, LLM-based retrieval 128, 163, knowledge graph enhancement techniques 108, 237, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 199, 177, 215.
-We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 97, 94, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes.
-Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 165, 18, 73 and open-source models 7, 204, 156, 44 alongside domain adaptation techniques 58, 19 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
+Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 24, 25 and dense retrieval methods 26, 27, LLM-based retrieval 4, 28, knowledge graph enhancement techniques 29, 30, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 31, 19, 32.
+We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 33, 34, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes.
+Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 35, 36, 37 and open-source models 38, 39, 40, 41 alongside domain adaptation techniques 6, 42 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
 Retriever
-Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 65.
-In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 218, 203.
-The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains190.
+Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 11.
+In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 43, 44.
+The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains45.
 Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation.
@@
 Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching.
-These methods, exemplified by algorithms like BM25 (Best Matching 25) 182 and TF-IDF (Term Frequency-Inverse Document Frequency) 106, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning.
+These methods, exemplified by algorithms like BM25 (Best Matching 25) 46 and TF-IDF (Term Frequency-Inverse Document Frequency) 47, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning.
 Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
 Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems.
-These papers demonstrate various applications of sparse retrieval, from pure BM25 42, 71 implementations to hybrid approaches 120, 13, 3, 115 that combine sparse and dense retrieval techniques.
-Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 109.
-Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 234, 88, 92.
+These papers demonstrate various applications of sparse retrieval, from pure BM25 48, 49 implementations to hybrid approaches 50, 51, 52, 53 that combine sparse and dense retrieval techniques.
+Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 54.
+Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 55, 56, 57.
 Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
 In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems.
-Established search tools including ElasticSearch 51, Whoosh 206, and Entrez API 56 also prove popular, offering accessible implementation options without requiring extensive customization.
-A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 231, 120, 13, 3, 115, leveraging both lexical precision and semantic understanding.
+Established search tools including ElasticSearch 58, Whoosh 59, and Entrez API 60 also prove popular, offering accessible implementation options without requiring extensive customization.
+A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 61, 50, 51, 52, 53, leveraging both lexical precision and semantic understanding.
 The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
@@
 Method Name Retrieval Method Task Year
-Clinfo.ai 141 Entrez API Medical QA 2023
-MEDRAG toolkit 231 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
-SeRTS 88 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
-Dual RAG System 120 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
-MedExpQA 13 BM25 and MedCPT Medical Question Answering 2024
-CliniqIR 3 BM25 and MedCPT Diagnostic decision support 2024
-VAIV Bio-Discovery 115 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
-Genomics analysis GPT 142 Azure AI Search with keyword matching Genetic variant annotation 2025
-Customized GPT 71 BM25 algorithm Medical QA 2025
-Two-Layer RAG 42 Whoosh (Okapi BM25F ranking) Medical question answering 2025
-LITURAt 169 Entrez API for PubMed retrieval Scientific data analysis 2025
+Clinfo.ai 14 Entrez API Medical QA 2023
+MEDRAG toolkit 61 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
+SeRTS 56 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
+Dual RAG System 50 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
+MedExpQA 51 BM25 and MedCPT Medical Question Answering 2024
+CliniqIR 52 BM25 and MedCPT Diagnostic decision support 2024
+VAIV Bio-Discovery 53 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
+Genomics analysis GPT 18 Azure AI Search with keyword matching Genetic variant annotation 2025
+Customized GPT 49 BM25 algorithm Medical QA 2025
+Two-Layer RAG 48 Whoosh (Okapi BM25F ranking) Medical question answering 2025
+LITURAt 62 Entrez API for PubMed retrieval Scientific data analysis 2025
 Dense Retrieval
-Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 246.
-These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 190.
+Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 63.
+These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 45.
 For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval.
@@
 There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems.
-As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 183, and Sentence Transformer 207 have been widely adopted in recent biomedical retrieval systems.
+As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 64, and Sentence Transformer 65 have been widely adopted in recent biomedical retrieval systems.
 The continued popularity of these models is evident from numerous recent implementations.
-Additionally, newer embedding models such as BGE-embedding 228, Dense Passage Retriever 111 and GIST Large Embedding 193 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks.
-Several other widely-used PLMs, including FastText 26, RoBERTa 137, and SimCSE 63, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
+Additionally, newer embedding models such as BGE-embedding 66, Dense Passage Retriever 67 and GIST Large Embedding 68 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks.
+Several other widely-used PLMs, including FastText 69, RoBERTa 70, and SimCSE 71, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
 Table 2: Classification of dense retriever types in biomedical RAG systems
@@
 Type 1: PLMs
-AskFDALabel 19 Sentence Transformer 207 FDA drug labeling extraction 2024
-CaLM 168 BGE-embedding 228 with Chroma DB Supporting caregivers FM 2024
-RAG-HPO 67 FastEmbed 205 Rare genetic disorders automated deep phenotyping 2024
-RALL 76 Dense Passage Retriever 111 Lay language generation 2024
-RT 129 BERT embeddings 46 Few-shot medical NER 2024
-RAG-GPT 233 BGE-embedding 228 Breast cancer nursing care QA 2024
-GNQA 41 HNSW graphs 147 Medical QA 2024
-Klang et al. 116 GIST Large Embedding 193 with FAISS ICD-10-CM coding 2024
-RAGPR 248 SBERT embeddings 183 Personalized physician recommendations 2025
-JGCLLM 61 GLuCoSE-base-ja vectors 60 Genetic counseling support 2025
+AskFDALabel 42 Sentence Transformer 65 FDA drug labeling extraction 2024
+CaLM 26 BGE-embedding 66 with Chroma DB Supporting caregivers FM 2024
+RAG-HPO 72 FastEmbed 73 Rare genetic disorders automated deep phenotyping 2024
+RALL 27 Dense Passage Retriever 67 Lay language generation 2024
+RT 74 BERT embeddings 75 Few-shot medical NER 2024
+RAG-GPT 76 BGE-embedding 66 Breast cancer nursing care QA 2024
+GNQA 77 HNSW graphs 78 Medical QA 2024
+Klang et al. 79 GIST Large Embedding 68 with FAISS ICD-10-CM coding 2024
+RAGPR 80 SBERT embeddings 64 Personalized physician recommendations 2025
+JGCLLM 81 GLuCoSE-base-ja vectors 82 Genetic counseling support 2025
 Type 2: Commercial APIs
-DRAGON-AI 209 OpenAI Text-Embedding 164 with Chroma DB Ontology generation 2024
-RISE 217 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
-Dual RAG 121 UPstage API and OpenAI Text-Embedding Diabetes management 2024
-FAVOR-GPT 130 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
-Endo-chat 59 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
-ChatENT 139 Open AI Text-Embedding Medical QA in otolaryngology 2024
-GuideGPT 198 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
-RECTIFIER 214 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
-RAP 20 Amazon Titan Text Embeddings v2 14 Nutrition-related question answering 2025
-InfectA-Chat 186 Open AI Text-Embedding Infectious disease monitoring and QA 2025
-DCRAG 107 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
+DRAGON-AI 83 OpenAI Text-Embedding 84 with Chroma DB Ontology generation 2024
+RISE 85 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
+Dual RAG 86 UPstage API and OpenAI Text-Embedding Diabetes management 2024
+FAVOR-GPT 87 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
+Endo-chat 88 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
+ChatENT 89 Open AI Text-Embedding Medical QA in otolaryngology 2024
+GuideGPT 90 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
+RECTIFIER 91 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
+RAP 92 Amazon Titan Text Embeddings v2 93 Nutrition-related question answering 2025
+InfectA-Chat 94 Open AI Text-Embedding Infectious disease monitoring and QA 2025
+DCRAG 95 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
 Type 3: Domain adaptation PLMs
-PM-Search 99 BioBERT 123 Clinical literature retrieval 2022
-LADER 98 PubMedBERT 75 Biomedical literature retrieval 2023
-CLEAR 140 BioBERT Clinical information extraction 2025
-KG-RAG 194 PubMedBERT Biomedical Multiple-choice questions 2024
-WeiseEule 6 MedCPT, BioBERT, BioGPT 145 with Pinecone DB Biomedical QA 2024
+PM-Search 24 BioBERT 96 Clinical literature retrieval 2022
+LADER 97 PubMedBERT 98 Biomedical literature retrieval 2023
+CLEAR 99 BioBERT Clinical information extraction 2025
+KG-RAG 100 PubMedBERT Biomedical Multiple-choice questions 2024
+WeiseEule 101 MedCPT, BioBERT, BioGPT 102 with Pinecone DB Biomedical QA 2024
 Type 4: LLM-based embedding
-BiomedRAG 128 MedLLaMA 13b Biomedical NLP tasks 2024
-SurgeryLLM 163 GPT4All 16 with Chroma DB Surgical decision support 2024
-Myers et al. 160 LLM2Vec (for comparison) 22 Clinical Information retrieval 2024
+BiomedRAG 4 MedLLaMA 13b Biomedical NLP tasks 2024
+SurgeryLLM 28 GPT4All 103 with Chroma DB Surgical decision support 2024
+Myers et al. 104 LLM2Vec (for comparison) 105 Clinical Information retrieval 2024
 Type 5: Domain adaptation retriever
-Self-BioRAG 93 MedCPT 97 Medical QA 2024
-WeiseEule 6 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
-CliniqIR 3 MedCPT and BM25 Diagnostic decision support 2024
-MEDRAG 6 BM25, SPECTER 37, Contriever 92and MedCPT Biomedical RAG Tool 2024
-RAMIE 242 MedCPT, Contriever and BMRetriever 234 Biomedical IR about dietary supplements 2025
+Self-BioRAG 106 MedCPT 33 Medical QA 2024
+WeiseEule 101 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
+CliniqIR 52 MedCPT and BM25 Diagnostic decision support 2024
+MEDRAG 101 BM25, SPECTER 22, Contriever 57and MedCPT Biomedical RAG Tool 2024
+RAMIE 107 MedCPT, Contriever and BMRetriever 55 Biomedical IR about dietary supplements 2025
 Type 2: Commercial embedding APIs.
@@
 While OpenAI’s offerings dominate, alternatives are emerging.
-The RAP system 20 utilizes Amazon’s Titan Text Embeddings v2 14 for nutrition-related question answering.
-Other commercial APIs such as Anthropic’s Voyage-2 17 and Google’s Vertex AI embeddings 74, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
+The RAP system 92 utilizes Amazon’s Titan Text Embeddings v2 93 for nutrition-related question answering.
+Other commercial APIs such as Anthropic’s Voyage-2 108 and Google’s Vertex AI embeddings 109, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
 Type 3: Domain adaptation PLMs.
-Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 123, 75.
-These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 146.
-By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 146.
-As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 75, BioBERT 123 and BioGPT 145, which have been widely adopted in biomedical RAG systems.
-For example, PM-Search 99 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision.
-Similarly, PubMedBERT has been successfully deployed in LADER 98 for biomedical literature retrieval and in KG-RAG 194 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks.
-More recent systems like WeiseEule 6 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 97 to achieve enhanced performance in biomedical QA.
+Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 96, 98.
+These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 110.
+By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 110.
+As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 98, BioBERT 96 and BioGPT 102, which have been widely adopted in biomedical RAG systems.
+For example, PM-Search 24 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision.
+Similarly, PubMedBERT has been successfully deployed in LADER 97 for biomedical literature retrieval and in KG-RAG 100 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks.
+More recent systems like WeiseEule 101 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 33 to achieve enhanced performance in biomedical QA.
 Type 4: LLM-based embedding.
 Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context.
-With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 247.
+With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 12.
 Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches.
-As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 128, which utilizes MedLLaMA 13b 224, SurgeryLLM 163, which employs GPT4All 16, and the comparative study by Myers et al.
-160, which evaluates LLM2Vec 22 for clinical information retrieval applications.
+As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 4, which utilizes MedLLaMA 13b 111, SurgeryLLM 28, which employs GPT4All 103, and the comparative study by Myers et al.
+104, which evaluates LLM2Vec 105 for clinical information retrieval applications.
 However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations.
-Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 138, Landmark Embedding 143, which provide promising dense retrieval alternatives in biomedical RAG.
+Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 112, Landmark Embedding 113, which provide promising dense retrieval alternatives in biomedical RAG.
 Type 5: Domain adaptation retrievers.
@@
 Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed.
-As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 92, MedCPT 97, SPECTER 37, BMRetriever 234, and the MEDRAG toolkit 128.
-These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 93, 3, 242.
+As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 57, MedCPT 33, SPECTER 22, BMRetriever 55, and the MEDRAG toolkit 4.
+These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 106, 52, 107.
 For example, Jeong et al.
 106 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach.
-Similarly, the MEDRAG framework 6 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool.
-RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 242.
+Similarly, the MEDRAG framework 101 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool.
+RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 107.
 These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
 Dense Retriever Infrastructure.
-Vector stores provide an infrastructure element for dense retrievers in semantic search systems 80.
+Vector stores provide an infrastructure element for dense retrievers in semantic search systems 114.
 These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers.
@@
 Store Description
-Chroma 36 AI-native open-source vector database with built-in functionality
-Faiss 52 Library for efficient similarity search and clustering of dense vectors
-Pinecone 172 Production-ready vector database for similarity search at scale
-Weaviate 220 Graph-based vector search engine with semantic capabilities
-HNSW 147 Hierarchical navigable small world graphs for approximate nearest neighbor search
+Chroma 115 AI-native open-source vector database with built-in functionality
+Faiss 116 Library for efficient similarity search and clustering of dense vectors
+Pinecone 117 Production-ready vector database for similarity search at scale
+Weaviate 118 Graph-based vector search engine with semantic capabilities
+HNSW 78 Hierarchical navigable small world graphs for approximate nearest neighbor search
 Knowledge Graph Enhanced Retrieval
 Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods.
-Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability171.
+Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability119.
 Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source.
-As we can notice, the application areas vary significantly, including medical question answering 194, knowledge extraction 47, and diagnosis prediction 64.
+As we can notice, the application areas vary significantly, including medical question answering 100, knowledge extraction 120, and diagnosis prediction 17.
 Table 4: Recent biomedical RAG studies based on knowledge graphs
 Method Task Knowledge Graph Used Year
-KG-RAG 194 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
-HEALIE 108 Personalized medical content generation HEALIE KG (customized) 2024
-KRAGEN 152 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
-Ascle 237 Various medical text generation tasks UMLS 2024
-DALK 125 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
-Gilbert et al. 69 Medical information structuring and interlinking Not specified 2024
-NetMe 2.0 47 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
-NEKO 229 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
-DR.KNOWS 64 Diagnosis prediction from EHRs UMLS 2025
-ESCARGOT 151 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
+KG-RAG 100 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
+HEALIE 29 Personalized medical content generation HEALIE KG (customized) 2024
+KRAGEN 121 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
+Ascle 30 Various medical text generation tasks UMLS 2024
+DALK 122 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
+Gilbert et al. 123 Medical information structuring and interlinking Not specified 2024
+NetMe 2.0 120 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
+NEKO 124 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
+DR.KNOWS 17 Diagnosis prediction from EHRs UMLS 2025
+ESCARGOT 125 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
 As shown in Fig 2, the integration of knowledge graphs with LLMs in biomedical RAG applications can be analyzed through four key components:
+Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
   • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories.
-NetMe 2.0 47 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 229 mines PubMed literature for synthetic biology knowledge graphs.
-Differently, KG-RAG 194 utilizes the pre-existing SPOKE biomedical knowledge graph.
-  • Query to knowledge graph transformation encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph.
-ESCARGOT 151 employs Cypher queries for precise retrieval, KRAGEN 152 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 64 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
+NetMe 2.0 120 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 124 mines PubMed literature for synthetic biology knowledge graphs.
+Differently, KG-RAG 100 utilizes the pre-existing SPOKE biomedical knowledge graph.
+  • Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph.
+ESCARGOT 125 employs Cypher queries for precise retrieval, KRAGEN 121 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 17 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
   • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query.
-DR.KNOWS 64 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 152 applies graph-of-thoughts for multi-hop relationships, and HEALIE 108 extracts paths incorporating both medical knowledge and socio-economic factors.
+DR.KNOWS 17 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 121 applies graph-of-thoughts for multi-hop relationships, and HEALIE 29 extracts paths incorporating both medical knowledge and socio-economic factors.
   • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency.
-KG-RAG 194 implements token-optimized retrieval to reduce consumption, ESCARGOT 151 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 47 transforms graph data into natural language prompts via its Graph-RAG module.
-Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
+KG-RAG 100 implements token-optimized retrieval to reduce consumption, ESCARGOT 125 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 120 transforms graph data into natural language prompts via its Graph-RAG module.
 Multi-modal Retrieval
 Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements.
-These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 4.
-As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 215, prescription label interpretation 208, lung cancer staging 211, digital pathology analysis 162, and automated radiology report generation 178, 199, 225, 226, 238.
+These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 126.
+As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 32, prescription label interpretation 127, lung cancer staging 128, digital pathology analysis 129, and automated radiology report generation 130, 31, 131, 132, 133.
 Figure 3: Architectural workflow of multimodal biomedical RAG system
@@
 Model Task Modalities Year
-Ranjit et al. 178 Chest X-Ray report generation Chest X-ray and text 2023
-FactMM-RAG 199 Radiology report generation X-Ray and text 2024
-MMed-RAG 225 Medical VQA & Report generation Medical images and text 2024
-MEAG 215 Amblyopia diagnosis Eye tracking and text 2024
-Tozuka et al.  211 Lung cancer staging CT findings and TNM classifications 2024
-GPT4DFCI-RAG 162 Digital pathology QA Pathology images and text 2024
-Raminedi et al. 177 Radiology report generation X-ray images and text 2024
-RULE 226 Medical VQA & Report generation Medical images and text 2024
-Thetbanthad et al. 208 Prescription label identification Image (labels) and text 2025
-Hu et al. 87 Pathology report generation Whole slide images and text 2025
-STREAM 238 Chest X-ray report generation Chest X-ray and text 2025
+Ranjit et al. 130 Chest X-Ray report generation Chest X-ray and text 2023
+FactMM-RAG 31 Radiology report generation X-Ray and text 2024
+MMed-RAG 131 Medical VQA & Report generation Medical images and text 2024
+MEAG 32 Amblyopia diagnosis Eye tracking and text 2024
+Tozuka et al.  128 Lung cancer staging CT findings and TNM classifications 2024
+GPT4DFCI-RAG 129 Digital pathology QA Pathology images and text 2024
+Raminedi et al. 19 Radiology report generation X-ray images and text 2024
+RULE 132 Medical VQA & Report generation Medical images and text 2024
+Thetbanthad et al. 127 Prescription label identification Image (labels) and text 2025
+Hu et al. 134 Pathology report generation Whole slide images and text 2025
+STREAM 133 Chest X-ray report generation Chest X-ray and text 2025
 Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system.
 The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture.
-While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 4.
-The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 32 for images, specialized encoders for time-series data 32); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge.
-For instance, Ranjit et al.178, FactMM-RAG 199 and STREAM 238 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report.
-MMed-RAG 225 and RULE 226 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation.
-The MEAG 215 system uses specialized encoders for eye movement data alongside text, Raminedi et al.
+While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 126.
+The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 135 for images, specialized encoders for time-series data 135); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge.
+For instance, Ranjit et al.130, FactMM-RAG 31 and STREAM 133 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report.
+MMed-RAG 131 and RULE 132 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation.
+The MEAG 32 system uses specialized encoders for eye movement data alongside text, Raminedi et al.
 129 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al.
@@
 Reranker
-Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 190.
+Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 45.
 Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component.
@@
 Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies.
-Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 3 and MEDRAG 6 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 209 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity.
-Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 6 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval.
-Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 94 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 5 utilizes a specialized cross-encoder model known as BGE M3-Embedding 31 to enhance clinical document relevance scoring.
-Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 97 trains a specific reranker denoted as CrossEnc.
-Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 41 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 39 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7.
+Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 52 and MEDRAG 101 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 83 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity.
+Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 101 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval.
+Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 34 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 136 utilizes a specialized cross-encoder model known as BGE M3-Embedding 137 to enhance clinical document relevance scoring.
+Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 33 trains a specific reranker denoted as CrossEnc.
+Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 77 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 138 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7.
 (on a 10 point scale).
 Finally, hybrid approaches combine multiple strategies to enhance performance.
-For example, Clinfo.ai 141 combines LLM classification with BM25 to re-rank retrieved articles.
+For example, Clinfo.ai 14 combines LLM classification with BM25 to re-rank retrieved articles.
 Generation
 The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries.
-In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 122.
+In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 139.
 This section examines various generation models employed in recent biomedical RAG studies.
@@
 As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception.
-This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 247.
-Among open-source LLMs, LLaMa 2 210 is the most popular, with 15 biomedical applications built upon it.
-LLaMa 3 7, Mistral 7B 155, and T5 176 are also widely used.
+This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 12.
+Among open-source LLMs, LLaMa 2 140 is the most popular, with 15 biomedical applications built upon it.
+LLaMa 3 38, Mistral 7B 141, and T5 142 are also widely used.
 These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems.
-Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 156 and Gemma 3 72.
+Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 40 and Gemma 3 143.
 Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
 For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance.
-Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 217, 84, 119, thereby offering viable alternatives to ChatGPT.
+Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 85, 144, 15, thereby offering viable alternatives to ChatGPT.
 Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
 In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications.
-Recent studies 166, 242, 11 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance.
+Recent studies 145, 107, 146 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance.
 Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
@@
 Open-Source LLMs
-T5 176 0.06-11B Encoder-Decoder VAIV 115, CLEAR 140, DR.KNOWS 64
-ChatGLM3 70 6B Decoder-only COPD 244, ChatZOC 144
-LLaMA 2 210 7-70B Decoder-only Guo et al.
-76, KREIMEYER et al.
-117, CaLM 168, MMRAG 241, Yu et al.
-239, RAMIE 242, EyeGPT 33, KG-RAG 194, JMLR 219, BiomedRAG 128, Self-BioRAG 93, CaLM 168, Du et al.
-49, SurgeryLLM 163, Kreimeyer et al.
-117
-LLaMA 3 7 8-70B Decoder-only MMRAG 241, Fatharanihttps et al. 54, RAMIE 242, Woo et al. 222, RAG-HPO 67, i-MedRAG 232, Hewitt et al. 84, PodGPT 94, SurgeryLLM 163
-Phi-3 Mini 154 3.8B Decoder-only Fatharanihttps et al. 54
-Mistral 7B 155 7B Decoder-only Boulos et al. 110, RAMIE 242, Woo et al. 222, RAGPR 248, LITURAt 169, Kreimeyer et al. 117
-QWen 175 32B Decoder-only NEKO 229, Thetbanthad 208
-Baichuan 216 7-13B Decoder-only ChatZOC 144
-Falcon 12 7-180B Decoder-only AskFDALabel 19, CaLM 168
-Zephyr 213 7B Decoder-only Moser et al. 159
-AceGPT 89 7-13B Decoder-only InfectA-Chat 186
-Gemma 62 1-27B Decoder-only PodGPT 94
-Deepseek R1 44 671B Decoder-only Feng et al. 55
+T5 142 0.06-11B Encoder-Decoder VAIV 53, CLEAR 99, DR.KNOWS 17
+ChatGLM3 147 6B Decoder-only COPD 148, ChatZOC 149
+LLaMA 2 140 7-70B Decoder-only Guo et al.
+27, KREIMEYER et al.
+150, CaLM 26, MMRAG 151, Yu et al.
+152, RAMIE 107, EyeGPT 153, KG-RAG 100, JMLR 3, BiomedRAG 4, Self-BioRAG 106, CaLM 26, Du et al.
+154, SurgeryLLM 28, Kreimeyer et al.
+150
+LLaMA 3 38 8-70B Decoder-only MMRAG 151, Fatharanihttps et al. 155, RAMIE 107, Woo et al. 156, RAG-HPO 72, i-MedRAG 157, Hewitt et al. 144, PodGPT 34, SurgeryLLM 28
+Phi-3 Mini 158 3.8B Decoder-only Fatharanihttps et al. 155
+Mistral 7B 141 7B Decoder-only Boulos et al. 159, RAMIE 107, Woo et al. 156, RAGPR 80, LITURAt 62, Kreimeyer et al. 150
+QWen 160 32B Decoder-only NEKO 124, Thetbanthad 127
+Baichuan 161 7-13B Decoder-only ChatZOC 149
+Falcon 162 7-180B Decoder-only AskFDALabel 42, CaLM 26
+Zephyr 163 7B Decoder-only Moser et al. 164
+AceGPT 165 7-13B Decoder-only InfectA-Chat 94
+Gemma 166 1-27B Decoder-only PodGPT 34
+Deepseek R1 41 671B Decoder-only Feng et al. 167
 Commercial LLMs
-ChatGPT-3.5/4/4o 165 Proprietary Decoder-only Woo et al.
-222, RAGPR 248, ChatZOC 144, KG-RAG 194, Clinfo.ai 141, Yu et al.
-239, ChatENT 139, CaLM 168, Rau et al.
-180, Endo-chat 59, BiomedRAG 128, Lu 142, Tan 15, GNQA 40, Puts 173, DRAGON-AI 209, Choi 35, RISE 217, Du et al.
-49,i-MedRAG 232 , Hewitt et al.
-84, RECTIFIER 214, Kainer et al.
-107, Gong et al.
-71, VAIA 115, RareDxGPT 240, Selcuk et al.
-186, FAVOR-GPT 130, Markey et al.
-148,
-Claude-3.5 18 Proprietary Decoder-only Woo et al. 222, RISE 217, Hewitt et al. 84
-Gemini 73 Proprietary Multimodal Upadhyaya et al. 215, Tozuka et al. 211, MEREDITH 119
+ChatGPT-3.5/4/4o 35 Proprietary Decoder-only Woo et al.
+156, RAGPR 80, ChatZOC 149, KG-RAG 100, Clinfo.ai 14, Yu et al.
+152, ChatENT 89, CaLM 26, Rau et al.
+168, Endo-chat 88, BiomedRAG 4, Lu 18, Tan 169, GNQA 77, Puts 170, DRAGON-AI 83, Choi 171, RISE 85, Du et al.
+154,i-MedRAG 157 , Hewitt et al.
+144, RECTIFIER 91, Kainer et al.
+95, Gong et al.
+49, VAIA 53, RareDxGPT 16, Selcuk et al.
+94, FAVOR-GPT 87, Markey et al.
+172,
+Claude-3.5 36 Proprietary Decoder-only Woo et al. 156, RISE 85, Hewitt et al. 144
+Gemini 37 Proprietary Multimodal Upadhyaya et al. 32, Tozuka et al. 128, MEREDITH 15
 Biomedical Specialized LLMs
-MedAlpaca 79 7-13B LLaMA-based RAMIE 242
-PMC-LLaMA 223 7B LLaMA-based RAMIE 242, BiomedRAG 128, Ozaki et al. 166
-BioMistral 118 7B Mistral-based RAMIE 242, pRAGe 28
-MEDITRON 34 7-70B LLaMA-based Ozaki et al. 166, Alkaeed et al. 11
-Meerkat 113 7B Mistral-based Sohn et al. 192
+MedAlpaca 173 7-13B LLaMA-based RAMIE 107
+PMC-LLaMA 174 7B LLaMA-based RAMIE 107, BiomedRAG 4, Ozaki et al. 145
+BioMistral 175 7B Mistral-based RAMIE 107, pRAGe 176
+MEDITRON 177 7-70B LLaMA-based Ozaki et al. 145, Alkaeed et al. 146
+Meerkat 178 7B Mistral-based Sohn et al. 179
 Domain Adaptation
-Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 189.
+Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 180.
 This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models.
@@
 Retriever Stage Adaptation.
-Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 92 and MedCPT 97) enable the development of more semantically rich biomedical embeddings.
+Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 57 and MedCPT 33) enable the development of more semantically rich biomedical embeddings.
 These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships.
-Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 234) and domain-specific architectural modifications (exemplified by LADER 98 with PubMedBERT initialization).
+Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 55) and domain-specific architectural modifications (exemplified by LADER 97 with PubMedBERT initialization).
 These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
 Reranker Stage Adaptation.
-The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 97) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 128).
+The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 33) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 4).
 Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage.
@@
 LLM Stage Adaptation.
-At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 58, SeRTS 88), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 19), and continued pre-training (RALL 76, ChatENT 139) enables more accurate synthesis and explication of biomedical information.
-JLMR 219 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions.
+At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 6, SeRTS 56), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 42), and continued pre-training (RALL 27, ChatENT 89) enables more accurate synthesis and explication of biomedical information.
+JLMR 3 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions.
 These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
@@
 Knowledge Source Datasets
-The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 66, 53.
+The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 181, 1.
 These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content.
-High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 90, 139, 93.
+High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 182, 89, 106.
 Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology.
 To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format.
-This integration enables systems to comprehensively address the multifaceted nature of medical queries 6, 234.
+This integration enables systems to comprehensively address the multifaceted nature of medical queries 101, 55.
 Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies.
-Among these sources, PubMed, UMLS 25, PMC, and Medical textbooks 95 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections.
-Additionally, we notice that several studies 35, 49, 248 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 120, 150, 233, 20, 89, 187 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
+Among these sources, PubMed, UMLS 183, PMC, and Medical textbooks 184 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections.
+Additionally, we notice that several studies 171, 154, 80 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 50, 185, 76, 92, 165, 186 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
 Table 7: Comprehensive Medical Knowledge Sources
 Source Type Description Studies
-MIMIC-IV 103 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 159
-PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 141, MEDRAG 231, Self-BioRAG 93, JMLR 219, Ascle 237, RISE 217, MEREDITH 119, LITURAt 169, DR.KNOWS 64, Aftab 6, CliniqIR 3, PM-Search 99, Kreimeyer 117, Kim 115
-PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 93, Soong et al. 196, PodGPT 94
-TripClick 181 Literature Biomedical literature search and retrieval dataset LADER 98
-GNQA 40 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 40
-UMLS 25 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 128, CLEAR 140, Guo 76, CliniqIR 3
-ICD-10 30 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 173
-HPO 102 Knowledge Base Human phenotype ontology Albayrak et al. 9, RAG-HPO 67, Kainer 107
-MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 234
-AlzKB 151 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 151, KRAGEN 152
-DrugBank 221 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 234, Kim 115
-SPOKE 158 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 194
-COD 144 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 144
-HEALIE KG108 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 108
-ADRD Caregiving 168 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 168
-StatPearls 21 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 231, i-MedRAG 232
-Medical Textbooks 95 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 231, Self-BioRAG 93, JMLR 219, i-MedRAG 232
-Ophthalmology textbooks 33 Educational 14 specialized ophthalmology textbooks EyeGPT 33
-CheXpert 91 Image-Report 224,316 chest radiographs with associated reports CLEAR 140
+MIMIC-IV 187 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 164
+PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 14, MEDRAG 61, Self-BioRAG 106, JMLR 3, Ascle 30, RISE 85, MEREDITH 15, LITURAt 62, DR.KNOWS 17, Aftab 101, CliniqIR 52, PM-Search 24, Kreimeyer 150, Kim 53
+PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 106, Soong et al. 188, PodGPT 34
+TripClick 189 Literature Biomedical literature search and retrieval dataset LADER 97
+GNQA 77 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 77
+UMLS 183 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 4, CLEAR 99, Guo 27, CliniqIR 52
+ICD-10 190 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 170
+HPO 191 Knowledge Base Human phenotype ontology Albayrak et al. 192, RAG-HPO 72, Kainer 95
+MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 55
+AlzKB 125 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 125, KRAGEN 121
+DrugBank 193 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 55, Kim 53
+SPOKE 194 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 100
+COD 149 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 149
+HEALIE KG29 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 29
+ADRD Caregiving 26 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 26
+StatPearls 23 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 61, i-MedRAG 157
+Medical Textbooks 184 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 61, Self-BioRAG 106, JMLR 3, i-MedRAG 157
+Ophthalmology textbooks 153 Educational 14 specialized ophthalmology textbooks EyeGPT 153
+CheXpert 195 Image-Report 224,316 chest radiographs with associated reports CLEAR 99
 Medical Task Datasets
-Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 81, 135.
+Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 2, 10.
 These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
@@
 Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems.
-Apart from works that make use of public biomedical datasets, several research groups 61, 198, 217 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
+Apart from works that make use of public biomedical datasets, several research groups 81, 90, 85 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
 Table 8: Comprehensive Biomedical Tasks Datasets
 Dataset Year Task Studies
-ChemProt 202 2010 Information Extraction BiomedRAG 128, ChemProt
-DDI 185 2013 Information Extraction BiomedRAG 128, ChemProt
-NCBI 48 2014 Entity Recognition GeneGPT 100, NetMe 47, RT 129
-BioASQ 212 2015 Information Retrieval BMRETRIEVER 234, SeRTS 88
-MIMIC-III 105 2016 Text Summarization DR.KNOWS 64, CliniqIR 3
-BC5CDR 126 2016 Entity Recognition RT 129
-DC3 50 2019 Diagnostic classification CliniqIR 3
-MIMIC-CXR 104 2019 Text Summarization Ascle 237
-MedQuAD 23 2019 QA Ascle 237
-PubMedQA 96 2019 Multiple-choice BMRETRIEVER 234, Liévin et al. 133, PodGPT 94
-MMLU 82 2021 Multiple-choice Self-BioRAG 93, JMLR 219, i-MedRAG 232, InfectA-Chat 186, PodGPT 94
-MedQA 95 2021 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, i-MedRAG 232, PodGPT 94
-HealthSearchQA 188 2022 QA
-MedMCQA 167 2022 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, EyeGPT 33, PodGPT 94
-ChatDoctor 132 2023 Dialogue EyeGPT 33
-MedAlpaca 79 2023 Dialogue EyeGPT 33
-GeneTuring 86 2023 Genomics QA GeneGPT 100
-GIT 127 2023 Information Extraction BiomedRAG 128
-MIRAGE 231 2024 Multiple-choice & QA MEDRAG 231
-MedExpQA 13 2024 Multilingual Medical QA MedExpQA 13, PodGPT 94
-PubMedRS-200 141 2024 Medical QA Clinifo.ai 141
-MedicineQA 90 2024 Multi-round dialogue RagPULSE 90
-CELLS 76 2024 Lay language generation Guo 76
+ChemProt 196 2010 Information Extraction BiomedRAG 4, ChemProt
+DDI 197 2013 Information Extraction BiomedRAG 4, ChemProt
+NCBI 198 2014 Entity Recognition GeneGPT 199, NetMe 120, RT 74
+BioASQ 200 2015 Information Retrieval BMRETRIEVER 55, SeRTS 56
+MIMIC-III 20 2016 Text Summarization DR.KNOWS 17, CliniqIR 52
+BC5CDR 201 2016 Entity Recognition RT 74
+DC3 202 2019 Diagnostic classification CliniqIR 52
+MIMIC-CXR 203 2019 Text Summarization Ascle 30
+MedQuAD 204 2019 QA Ascle 30
+PubMedQA 205 2019 Multiple-choice BMRETRIEVER 55, Liévin et al. 206, PodGPT 34
+MMLU 207 2021 Multiple-choice Self-BioRAG 106, JMLR 3, i-MedRAG 157, InfectA-Chat 94, PodGPT 34
+MedQA 184 2021 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, i-MedRAG 157, PodGPT 34
+HealthSearchQA 208 2022 QA
+MedMCQA 209 2022 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, EyeGPT 153, PodGPT 34
+ChatDoctor 210 2023 Dialogue EyeGPT 153
+MedAlpaca 173 2023 Dialogue EyeGPT 153
+GeneTuring 211 2023 Genomics QA GeneGPT 199
+GIT 212 2023 Information Extraction BiomedRAG 4
+MIRAGE 61 2024 Multiple-choice & QA MEDRAG 61
+MedExpQA 51 2024 Multilingual Medical QA MedExpQA 51, PodGPT 34
+PubMedRS-200 14 2024 Medical QA Clinifo.ai 14
+MedicineQA 182 2024 Multi-round dialogue RagPULSE 182
+CELLS 27 2024 Lay language generation Guo 27
 Applications
 The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval.
-This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 136.
+This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 213.
 In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
@@
 Medical Question Answering
-Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 194.
+Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 100.
 This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
 Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval.
-Clinfo.ai 141 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence.
-RALL 76 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information.
-Self-BioRAG 93 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 6 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties.
+Clinfo.ai 14 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence.
+RALL 27 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information.
+Self-BioRAG 106 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 101 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties.
 These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
 Diagnostic and Treatment Decision Support
-Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 201.
+Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 214.
 They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
 Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties.
-Expert-Guided LLMs 119 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data.
-DR.KNOWS 64 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments.
-CliniqIR 3 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 163 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data.
-RECTIFIER 214 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data.
+Expert-Guided LLMs 15 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data.
+DR.KNOWS 17 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments.
+CliniqIR 52 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 28 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data.
+RECTIFIER 91 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data.
 By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
 Rare Disease Identification and Management
-Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 149.
+Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 215.
 RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
@@
 16 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases.
-RAG-HPO 67 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping.
+RAG-HPO 72 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping.
 Yang et al.
 216 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases.
-Some authors 114 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis.
+Some authors 217 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis.
 These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
 Clinical Report Generation
-Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 191.
-A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 78.
-This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 10.
-RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 230.
+Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 218.
+A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 219.
+This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 220.
+RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 7.
 Recent works have demonstrated the effectiveness of RAG in this domain.
-For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 179, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent.
-To enhance interpretability, some authors 8 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability.
-RAG pipelines 24 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors.
-Differently, LaB-RAG 195 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM.
-Lastly, some authors 200 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
+For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 221, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent.
+To enhance interpretability, some authors 222 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability.
+RAG pipelines 223 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors.
+Differently, LaB-RAG 224 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM.
+Lastly, some authors 225 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
 Precision Medicine Applications
@@
 Genomic medicine involves analyzing genetic data to inform clinical decision-making.
-The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 29.
+The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 226.
 Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval.
-FAVOR-GPT 130 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases.
+FAVOR-GPT 87 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases.
 Lu et al.
 18 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts.
-RAG-HPO 67 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 61 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information.
-By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 142 and phenotypic and genotypic analysis 67, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
+RAG-HPO 72 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 81 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information.
+By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 18 and phenotypic and genotypic analysis 72, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
 Personalized Health Management
-Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 174.
+Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 227.
 This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
 Recent research has produced several personalized health management systems employing different knowledge retrieval strategies.
-HEALIE 108 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients.
-RISE 217 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements.
-A Dual RAG System 120 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations.
-RAP 20 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions.
+HEALIE 29 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients.
+RISE 85 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements.
+A Dual RAG System 50 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations.
+RAP 92 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions.
 These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
 Healthcare Education
-Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 124.
+Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 228.
 These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
-ChatENT 139 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS).
-EyeGPT 33 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries.
+ChatENT 89 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS).
+EyeGPT 153 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries.
 Sevgi et al.
 186 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval.
-In nursing education, Zhao et al 245 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification.
+In nursing education, Zhao et al 229 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification.
 By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.
@@
 Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations.
-68.
+230.
 Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches.
-PodGPT 94 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development.
-LITURAt 169 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification.
+PodGPT 34 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development.
+LITURAt 62 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification.
 These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
@@
 Trustworthiness
-Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 53, 45.
+Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 1, 231.
 Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
 Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns.
-Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 66.
+Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 181.
 Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications.
@@
 A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources.
-For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 122.
+For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 139.
 This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems.
 Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process.
-Due to the uneven distribution of demographic attributes in training corpora 81, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases.
+Due to the uneven distribution of demographic attributes in training corpora 2, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases.
 Addressing the fairness concerns also represents an important direction for future research.
@@
 This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets.
-The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 161, 101.
+The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 232, 233.
 This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
@@
 Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms.
-Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 157, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
+Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 234, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
 Multimodality
 Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g.
-X-ray 104, 170, 134), time series (e.g.
-ECG 239), sound recordings (e.g.
-lung sounds recorded 57), and video 77.
+X-ray 203, 235, 236), time series (e.g.
+ECG 152), sound recordings (e.g.
+lung sounds recorded 237), and video 238.
 Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored.
 Multi-modal data presents several intractable challenges for retrieval systems.
-For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 249.
+For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 239.
 Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression.
@@
 The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios.
-Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 53.
+Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 1.
 However, these models typically require high-performance computing clusters to operate effectively.
-For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 43.
-Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 27, 85.
+For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 240.
+Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 241, 242.
 Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive.
 Several promising approaches have emerged to address these constraints.
-One solution involves developing smaller, distilled LLMs derived from their larger counterparts 235.
-Models such as QWQ 175 and Mistral Small 3.1 156 demonstrate competitive performance despite having significantly reduced parameter counts.
-Applying Portable LLMs 112 as generation models is also a promising method.
+One solution involves developing smaller, distilled LLMs derived from their larger counterparts 243.
+Models such as QWQ 160 and Mistral Small 3.1 40 demonstrate competitive performance despite having significantly reduced parameter counts.
+Applying Portable LLMs 244 as generation models is also a promising method.
 Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services.
@@
 Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes.
-During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 153.
-Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 38 and the Health Insurance Portability and Accountability Act 2.
+During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 245.
+Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 246 and the Health Insurance Portability and Accountability Act 247.
 Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development.

### v3

@@
 Introduction
-Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 1, 2.
+Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks Fan2024RAG, he2023survey.
 These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval.
-The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 3.
+The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedicine domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes Wang2024JMLR.
 However, the application of LLMs in the biomedical domain faces two critical challenges.
-First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 4.
+First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination li2025biomedrag.
 This is particularly concerning in medical applications, where accuracy is paramount for patient safety.
-Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 5.
+Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge XIAO2025102888.
 This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
-Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 6.
-By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 1, 2, 5.
+Biomedicine RAG Dataset and Evaluation § IV Technology § III Preliminary § II Introduction § I Applications § V Discussion § VI Conclusion § VII Reranker Retriever Generation Domain Adaptation Medical Task Datasets Knowledge Source Datasets Evaluation Metrics Precision Medicine Applications Clinical Report Generation Clinical Decision Support Systems Healthcare Education Clinical Research Summary of dataset and evaluation approaches Summary of technological approaches Summary of RAG applications in biomedicine Challenges and Future Direction Knowledge Graph Enhanced Retrieval Dense Retrieval Sparse Retrieval Multi-modal Retrieval Commercial LLMs Open-Source LLMs Biomedical Specialized LLMs Retriever Stage Adaptation Reranker Stage Adaptation LLM Stage Adaptation Diagnostic and Treatment Decision Support Medical Question Answering Rare Disease Identification and Management Genomic Medicine Personalized Health Management
+Figure 1: Structure of this paper
+Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges frisoni2024generate.
+By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) reduce hallucinations by grounding responses in verified medical literature, thereby enhancing the accuracy and reliability of medical information processingFan2024RAG, he2023survey, XIAO2025102888.
 The implementation of RAG in the biomedical domain presents unique challenges and opportunities.
-Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 7.
-Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 8, 9.
-Concurrent to our survey, several related surveys focused on biomedical-related LLMs.
-For example, Liu et al.
-10 review medical LLMs and He et al.
-2 discuss LLMs for healthcare.
-Similarly, Xiao 5 conducted a comprehensive overview of LLMs in medicine.
-Other discussed RAG systems in general domains, such as Fan et al.
-1 and Gao et al.
-11 focused on the architectures, training strategies, and application of RAG in general domains.
-Our work differs from existing surveys in concentrating on the application of RAG in the biomedical domain, systematically reviewing its components, datasets, and challenges.
+Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability xiong2024benchmarking.
+Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 10.1007/978-981-97-9437-9_2, santra2024regulatory.
+Therefore, a systematic survey of existing studies applying RAG in the biomedicine and healthcare domain is necessary to understand how these challenges are being addressed and to identify current research trends and gaps.
+We considered papers that are directly related to the application of RAG techniques in the biomedicine or healthcare domain.
+Eligible studies include those that incorporate RAG components—such as retrievers, rerankers(optional), and generation models—or apply RAG to improve model performance on biomedical tasks.
+Specifically, we included works applying RAG to biomedical named entity recognition (NER), biomedical question answering (QA), multiple-choice reasoning, diagnostic classification, and medical dialogue systems, including, but not limited to, these areas.
+Studies were required to demonstrate how RAG integration enhances the performance, accuracy, reliability, or factual grounding of biomedical models, systems, or applications.
+Studies that use RAG only as a baseline or comparative model without substantively integrating RAG into their methodological framework were excluded.
+To collect relevant literature, we systematically searched PubMed and Google Scholar for studies published between 2022 and 2025, ensuring coverage of the most recent developments in biomedical RAG research.
+In total, 143 papers met the inclusion criteria and were included in this survey.
+Concurrent to our work, several surveys have examined either biomedical LLMs or general-domain RAG systems.
+Surveys such as Liu et al.
+liu2024survey and He et al.
+he2023survey focus primarily on medical LLMs, emphasizing model evolution, training data, clinical applications, and safety considerations, but they do not analyze the RAG pipeline or retrieval–generation integration.
+Xiao et al.
+XIAO2025102888 further expand to multimodal LLMs in medicine, yet similarly concentrate on model architectures and clinical use cases rather than grounding mechanisms or retrieval-oriented methodologies.
+In contrast, recent surveys on RAG—such as Fan et al.
+Fan2024RAG and Gao et al.
+Gao2023RAGSurvey—provide systematic overviews of RAG architectures, training paradigms, and evaluation practices in general domains.
+However, they do not address the unique challenges of biomedical knowledge sources, medical-language characteristics, domain adaptation strategies, or multimodal medical data.
+Our survey bridges this gap by offering the first retrieval-centric and pipeline-level review of RAG techniques specifically tailored for biomedical applications.
+We synthesize advances from sparse/dense/LLM-based retrieval, biomedical rerankers, and domain-adapted generation models, and further integrate developments in knowledge graphs and multimodal retrieval.
+By unifying these threads, our work positions biomedical RAG as a distinct research area and provides a structured roadmap that helps developers and researchers understand current progress and effectively contribute to future advances.
+Table 1 provides a brief comparison of recent related survey.
+Table 1: Comparative summary of related survey papers
+Reference Year RAG Biomedical Pipeline Coverage Key Contributions
+He et al. he2023survey 2023 Generation only Provides an overview of medical LLM development, applications, and ethical considerations, but does not address retrieval or grounding mechanisms.
+Liu et al. liu2024survey 2024 Generation only Covers Med-LLM technologies, applications, and trustworthiness, yet lacks discussion of retrieval augmentation or biomedical evidence integration.
+Xiao et al. XIAO2025102888 2025 Generation only Reviews LLMs and MLLMs in medicine, emphasizing model architecture and clinical scenarios, but does not examine retrieval, reranking, or grounding.
+Fan et al.
+Fan2024RAG 2024 Retrieval + Generation Provides a structured overview of RAG architectures and training strategies in general domains, without biomedical datasets or domain-specific methods.
+Gao et al. Gao2023RAGSurvey 2023 Retrieval + Generation Introduces Naive, Advanced, and Modular RAG frameworks, but does not cover biomedical tasks, knowledge sources, or clinical evaluation.
+Ours 2025 Full pipeline: retrieval, reranking, generation, domain adaptation Offers the first domain-specific, pipeline-level survey of biomedical RAG, integrating sparse/dense/LLM-based retrievers, biomedical rerankers, domain-adapted generators, knowledge-graph RAG, multimodal retrieval, and medically grounded evaluation practices.
 Overall, this survey provides the following key contributions:
   1. 1.
-    A systematic analysis of RAG technologies specifically adapted for biomedical applications, including specialized retrieval methods, ranking algorithms, and generation models.
+    A comprehensive review of retrieval-augmented technologies tailored to the biomedical domain, covering sparse, dense, LLM-based, and domain-specialized retrievers, as well as biomedical rerankers and generation models.
   2. 2.
-    A detailed examination of medical datasets, considering both knowledge source datasets and medical task datasets.
+    A detailed examination of medical datasets and evaluation, considering both knowledge source datasets and medical task benchmarks, along with their evaluation methods.
   3. 3.
-    An exploration of current clinical applications and future directions, with particular attention to practical implementation challenges and emerging opportunities.
+    A practical, pipeline-oriented framework that examines existing biomedical RAG applications and offers developers and practitioners a clear roadmap for system implementation, while also highlighting future challenges and research directions.
+Fig.
+1 illustrates the overall structure of this survey.
+The remainder of the paper is organized as follows.
+Section II introduces the background and fundamental framework of biomedical RAG systems.
+Section III details the core technical components, including retrieval methods, ranking algorithms, and generation models.
+Section IV reviews the available datasets and evaluation methodologies.
+Section V discusses biomedical applications that leverage RAG.
+Section VI summarizes the key observations and discusses current challenges and future directions.
+Finally, Section VII concludes the paper.
+Preliminary
+To provide a clear conceptual and structural understanding of biomedical RAG systems, we first summarize their general workflow and key components.
+As Fig.
+2 shows, in a typical biomedical RAG workflow, the system can be formally characterized as a composition of sequential functions operating over a query and an external knowledge source.
+Given a user query $q$ and a heterogeneous knowledge collection $K$, the end-to-end RAG process generates a response $y$ through the functional pipeline:
+$y=G\big(q,\,S\big(q,\,R(q,K)\big)\big),$ (1)
+where each component $R$, $S$ and $G$ represents a distinct stage of the retrieval–generation paradigm.
+First, the Retriever $R(q,K)$ identifies a set of candidate documents relevant to the query and returns a document collection $D=\{d_{1},d_{2},\ldots,d_{n}\}\subseteq K$.
+Next, the Re-ranking Method $S(q,D)$ is an optional component of the RAG pipeline that refines the retrieved set by sorting and prioritizing documents according to task-specific relevance criteria, producing a re-ranked subset $D^{\prime}=[d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{m}]$ with $D^{\prime}\subseteq D$.
+The use of this component, as well as the choice of scoring strategy, varies across biomedical applications.
+Finally, the Generation Function $G(q,D^{\prime})$ synthesizes the retrieved evidence with the original query to produce the final output:
+$y=G(q,D^{\prime})=\text{LLM}\big(\text{prompt}(q,D^{\prime})\big),$ (2)
+where the prompting function integrates the query with the selected context, $\text{prompt}(q,D^{\prime})=q\oplus D^{\prime}$.
+This formulation describes RAG as a modular and extensible computational paradigm, whose behavior can be adapted to a wide range of biomedical tasks through different designs of the generation function $G$.
+To illustrate this flexibility, we formalize 4 representative instantiations of $G$ in biomedical applications:
+  1. 1.
+    Biomedical Text Classification.
+For classification-oriented tasks, the generation function $G$ selects the most probable label from the label set $\mathcal{C}$ under the model’s conditional distribution:
+    $G_{\text{cls}}(q,D^{\prime})=\arg\max_{c\in\mathcal{C}}\,p_{\theta}(c\mid\text{prompt}(q,D^{\prime})).$ (3)
+    This formulation encompasses a variety of biomedical classification scenarios, including diagnostic classification and multiple-choice question answering.
+  2. 2.
+    Biomedical Question Answering.
+For free-text biomedical QA, $G$ generates an answer sequence $y=[y_{1},\ldots,y_{T}]$ by sampling (or decoding) tokens from the conditional language model distribution:
+    $G_{\text{qa}}(q,D^{\prime})=\arg\max_{y}\,p_{\theta}(y\mid\text{prompt}(q,D^{\prime})),$ (4)
+    where $y$ represents the generated natural-language answer.
+  3. 3.
+    Clinical Dialogue or Consultation Systems. For multi-turn dialogue, $G$ conditions on the dialogue history $h$ in addition to the retrieved evidence:
+    $G_{\text{dialog}}(q,D^{\prime},h)=\arg\max_{y}\,p_{\theta}(y\mid h\oplus\text{prompt}(q,D^{\prime})),$ (5)
+    where $y$ is the next system response.
   4. 4.
-    A structured framework for understanding the interplay between retrieval mechanisms, medical knowledge sources, and language models in biomedical RAG systems.
-The remainder of this survey is organized as follows: Section 2 presents the background and the basic framework of biomedical RAG systems.
-Section 3 details the technical components, including retrieval methods, ranking algorithms, and generation models.
-Section 4 examines available datasets.
-Section 5 explores current clinical applications, and Section 6 discusses challenges and future directions.
-Background
-LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 12.
-Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 13.
-RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 11.
-This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 5.
-In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources.
-The re-ranking method then sorts these references based on their relevance.
-Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response.
-As Fig.
-1 shows, the basic framework of biomedical RAG consists of four key components 1: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
+    Biomedical Named Entity Recognition (NER). For sequence labeling tasks, $G$ predicts a sequence of token-level labels:
+    $G_{\text{ner}}(q,D^{\prime})=\big[\arg\max_{y_{t}\in\mathcal{L}}\,p_{\theta}(y_{t}\mid\text{prompt}(q,D^{\prime}),t)\big]_{t=1}^{T},$ (6)
+    where $\mathcal{L}$ is the label set and $T$ is the input sequence length.
+These formulations highlight how biomedical RAG systems form a unified retrieval–generation pipeline, while tailoring the generation function to meet the requirements of diverse downstream tasks.
+Beyond the task-specific instantiations of the generation function $G$, the RAG paradigm further benefits from the inherent extensibility of its external knowledge source.
+The knowledge repository $K$ is designed to be continuously updatable, enabling biomedical RAG systems to maintain timely access to newly emerging medical evidence.
+As new documents or knowledge collections—such as updated clinical guidelines, recently published biomedical literature, or domain-specific databases—become available, they can be seamlessly incorporated into the existing repository.
+Formally, the expanded knowledge source is expressed as $K_{\text{updated}}=K\cup K_{\text{new}}$, where $K_{\text{new}}$ denotes the newly added knowledge.
+This extensibility ensures that retrieval and generation remain grounded in the most current biomedical information, thereby enhancing the robustness and adaptability of RAG across diverse clinical and research applications.
 Figure 2: Basic biomedical RAG framework.
-RAG has been applied to a variety of biomedical applications.
-In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 14, enhances specialized decision-making through expert-guided LLMs 15, and facilitates rare disease diagnosis 16.
-This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 17 and advances genomic medicine through improved variant annotation and interpretation 18.
-In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 19.
-Across these applications, RAG leverages diverse data sources—including health records 20, 21, scientific literature 22, clinical guidelines 23, and genomic databases 18—to generate evidence-based responses that significantly improve both clinical care and research outcomes.
-The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
+RAG has been widely applied across diverse biomedical and healthcare scenarios, including clinical decision support, healthcare education, clinical report generation, and other related applications.
+Across these applications, RAG use diverse data sources-including health records johnson2016mimic, herrett2015cprd, scientific literature specter2020cohan, clinical guidelines Bashir2024, and genomic databases Lu2025BoostingGPT—to generate evidence-based responses that significantly improve both clinical care and research outcomes.
+A more detailed discussion of specific applications is provided in Section V.
 Technology
 In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information.
-Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 24, 25 and dense retrieval methods 26, 27, LLM-based retrieval 4, 28, knowledge graph enhancement techniques 29, 30, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 31, 19, 32.
-We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 33, 34, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes.
-Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 35, 36, 37 and open-source models 38, 39, 40, 41 alongside domain adaptation techniques 6, 42 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
+Our examination of retrieval mechanisms encompasses a range of approaches, including sparse Jin2022, zhang2024online and dense retrieval methods parmanto2024calm, Guo2024, LLM-based retrieval li2025biomedrag, ong2024surgeryllm, knowledge graph enhancement techniques Kakalou2024, Ascle, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts Sun2024FactAware, Raminedi2024MultimodalTransformer, Upadhyaya2024.
+We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents jin2023medcpt, jia2024podgpt, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes.
+Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source OpenAI2025TextGeneration, Anthropic2025GettingStarted, GoogleGemini2025 and open-source models MetaAI2024LLaMA3, MosaicML2023Introducing, mistral_small_24b_2024, deepseekr1_2024 alongside domain adaptation techniques frisoni2024generate, llm-regulatory-framework that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
 Retriever
-Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 11.
-In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 43, 44.
-The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains45.
+Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries Gao2023RAGSurvey.
+In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 10.1145/3648471, 10.1145/3462476.
+The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domainssivarajkumar2024clinical.
 Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation.
-In this section, we describe various retrieval methodologies.
-We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
+In this subsection, we present four main retrieval paradigms used in biomedical RAG systems.
+Sparse and dense retrieval represent the most commonly adopted approaches, whereas knowledge graph–enhanced and multimodal retrieval are designed to handle more complex data structures, such as structured knowledge graphs and multimodal biomedical information.
 Sparse Retrieval
 Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching.
-These methods, exemplified by algorithms like BM25 (Best Matching 25) 46 and TF-IDF (Term Frequency-Inverse Document Frequency) 47, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning.
+These methods, exemplified by algorithms like BM25 (Best Matching 25) bm25 and TF-IDF (Term Frequency-Inverse Document Frequency) Jones1972, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning.
 Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
 Table 2 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems.
-These papers demonstrate various applications of sparse retrieval, from pure BM25 48, 49 implementations to hybrid approaches 50, 51, 52, 53 that combine sparse and dense retrieval techniques.
-Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 54.
-Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 55, 56, 57.
+These papers demonstrate various applications of sparse retrieval, from pure BM25 Das2025, Gong2024 implementations to hybrid approaches Lee2024, ALONSO2024102938, Abdullahi2024, Kim2024 that combine sparse and dense retrieval techniques.
+Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization kamalloo:2024.
+Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements xu2024bmretriever, Hu2024, izacard2021contriever.
 Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
 In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems.
-Established search tools including ElasticSearch 58, Whoosh 59, and Entrez API 60 also prove popular, offering accessible implementation options without requiring extensive customization.
-A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 61, 50, 51, 52, 53, leveraging both lexical precision and semantic understanding.
+Established search tools including ElasticSearch ElasticsearchDocs, Whoosh WhooshDocs, and Entrez API EntrezAPI also prove popular, offering accessible implementation options without requiring extensive customization.
+A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques xiong-etal-2024-benchmarking, Lee2024, ALONSO2024102938, Abdullahi2024, Kim2024, leveraging both lexical precision and semantic understanding.
 The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
@@
 Method Name Retrieval Method Task Year
-Clinfo.ai 14 Entrez API Medical QA 2023
-MEDRAG toolkit 61 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
-SeRTS 56 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
-Dual RAG System 50 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
-MedExpQA 51 BM25 and MedCPT Medical Question Answering 2024
-CliniqIR 52 BM25 and MedCPT Diagnostic decision support 2024
-VAIV Bio-Discovery 53 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
-Genomics analysis GPT 18 Azure AI Search with keyword matching Genetic variant annotation 2025
-Customized GPT 49 BM25 algorithm Medical QA 2025
-Two-Layer RAG 48 Whoosh (Okapi BM25F ranking) Medical question answering 2025
-LITURAt 62 Entrez API for PubMed retrieval Scientific data analysis 2025
+Clinfo.ai Lozano2024ClinfoAI Entrez API Medical QA 2023
+MEDRAG toolkit xiong-etal-2024-benchmarking Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
+SeRTS Hu2024 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
+Dual RAG System Lee2024 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
+MedExpQA ALONSO2024102938 BM25 and MedCPT Medical Question Answering 2024
+CliniqIR Abdullahi2024 BM25 and MedCPT Diagnostic decision support 2024
+VAIV Bio-Discovery Kim2024 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
+Genomics analysis GPT Lu2025BoostingGPT Azure AI Search with keyword matching Genetic variant annotation 2025
+Customized GPT Gong2024 BM25 algorithm Medical QA 2025
+Two-Layer RAG Das2025 Whoosh (Okapi BM25F ranking) Medical question answering 2025
+LITURAt Peasley2025 Entrez API for PubMed retrieval Scientific data analysis 2025
 Dense Retrieval
-Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 63.
-These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 45.
+Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 10.1145/3637870.
+These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching sivarajkumar2024clinical.
 For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval.
 To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers.
-In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
-Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings.
+In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical RAG systems.
+Type 1: PLMs.
+PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings.
 There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems.
-As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 64, and Sentence Transformer 65 have been widely adopted in recent biomedical retrieval systems.
+As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT santander2022semantic, and Sentence Transformer thakur2020augmented have been widely adopted in recent biomedical retrieval systems.
 The continued popularity of these models is evident from numerous recent implementations.
-Additionally, newer embedding models such as BGE-embedding 66, Dense Passage Retriever 67 and GIST Large Embedding 68 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks.
-Several other widely-used PLMs, including FastText 69, RoBERTa 70, and SimCSE 71, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
+Additionally, newer embedding models such as BGE-embedding bge_embedding, Dense Passage Retriever Karpukhin2020 and GIST Large Embedding solatorio2024gistembed have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks.
+Several other widely-used PLMs, including FastText bojanowski2017enriching, RoBERTa liu2019roberta, and SimCSE gao-etal-2021-simcse, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
 Table 3: Classification of dense retriever types in biomedical RAG systems
@@
 Type 1: PLMs
-AskFDALabel 42 Sentence Transformer 65 FDA drug labeling extraction 2024
-CaLM 26 BGE-embedding 66 with Chroma DB Supporting caregivers FM 2024
-RAG-HPO 72 FastEmbed 73 Rare genetic disorders automated deep phenotyping 2024
-RALL 27 Dense Passage Retriever 67 Lay language generation 2024
-RT 74 BERT embeddings 75 Few-shot medical NER 2024
-RAG-GPT 76 BGE-embedding 66 Breast cancer nursing care QA 2024
-GNQA 77 HNSW graphs 78 Medical QA 2024
-Klang et al. 79 GIST Large Embedding 68 with FAISS ICD-10-CM coding 2024
-RAGPR 80 SBERT embeddings 64 Personalized physician recommendations 2025
-JGCLLM 81 GLuCoSE-base-ja vectors 82 Genetic counseling support 2025
+AskFDALabel llm-regulatory-framework Sentence Transformer thakur2020augmented FDA drug labeling extraction 2024
+CaLM parmanto2024calm BGE-embedding bge_embedding with Chroma DB Supporting caregivers FM 2024
+RAG-HPO garcia2024deepphenotyping FastEmbed fastembed Rare genetic disorders automated deep phenotyping 2024
+RALL Guo2024 Dense Passage Retriever Karpukhin2020 Lay language generation 2024
+RT li2024rt BERT embeddings devlin2019bert Few-shot medical NER 2024
+RAG-GPT xu2024evaluation BGE-embedding bge_embedding Breast cancer nursing care QA 2024
+GNQA darnell2024biomedical HNSW graphs malkov2016efficient Medical QA 2024
+Klang et al. klang2024assessing GIST Large Embedding solatorio2024gistembed with FAISS ICD-10-CM coding 2024
+RAGPR zheng2025integrating SBERT embeddings santander2022semantic Personalized physician recommendations 2025
+JGCLLM Fukushima2025 GLuCoSE-base-ja vectors Fukuchi2023 Genetic counseling support 2025
 Type 2: Commercial APIs
-DRAGON-AI 83 OpenAI Text-Embedding 84 with Chroma DB Ontology generation 2024
-RISE 85 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
-Dual RAG 86 UPstage API and OpenAI Text-Embedding Diabetes management 2024
-FAVOR-GPT 87 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
-Endo-chat 88 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
-ChatENT 89 Open AI Text-Embedding Medical QA in otolaryngology 2024
-GuideGPT 90 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
-RECTIFIER 91 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
-RAP 92 Amazon Titan Text Embeddings v2 93 Nutrition-related question answering 2025
-InfectA-Chat 94 Open AI Text-Embedding Infectious disease monitoring and QA 2025
-DCRAG 95 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
+DRAGON-AI toro2024dragonai OpenAI Text-Embedding openai_embeddings with Chroma DB Ontology generation 2024
+RISE wang2024diabetesrag OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
+Dual RAG lee2024dualrag UPstage API and OpenAI Text-Embedding Diabetes management 2024
+FAVOR-GPT li2024favorgpt Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
+Endo-chat fu2024application Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
+ChatENT long2024chatent Open AI Text-Embedding Medical QA in otolaryngology 2024
+GuideGPT steybe2025evaluation Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
+RECTIFIER unlu2024retrieval Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
+RAP azimi2025evaluation Amazon Titan Text Embeddings v2 aws_titan_embedding Nutrition-related question answering 2025
+InfectA-Chat selcuk2025infecta Open AI Text-Embedding Infectious disease monitoring and QA 2025
+DCRAG kainer2025effectiveness OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
 Type 3: Domain adaptation PLMs
-PM-Search 24 BioBERT 96 Clinical literature retrieval 2022
-LADER 97 PubMedBERT 98 Biomedical literature retrieval 2023
-CLEAR 99 BioBERT Clinical information extraction 2025
-KG-RAG 100 PubMedBERT Biomedical Multiple-choice questions 2024
-WeiseEule 101 MedCPT, BioBERT, BioGPT 102 with Pinecone DB Biomedical QA 2024
+PM-Search Jin2022 BioBERT 10.1093/bioinformatics/btz682 Clinical literature retrieval 2022
+LADER LADER PubMedBERT pubmedbert Biomedical literature retrieval 2023
+CLEAR lopez2025clinical BioBERT Clinical information extraction 2025
+KG-RAG Soman2024 PubMedBERT Biomedical Multiple-choice questions 2024
+WeiseEule Aftab2024 MedCPT, BioBERT, BioGPT luo2022biogpt with Pinecone DB Biomedical QA 2024
 Type 4: LLM-based embedding
-BiomedRAG 4 MedLLaMA 13b Biomedical NLP tasks 2024
-SurgeryLLM 28 GPT4All 103 with Chroma DB Surgical decision support 2024
-Myers et al. 104 LLM2Vec (for comparison) 105 Clinical Information retrieval 2024
+BiomedRAG li2025biomedrag MedLLaMA 13b Biomedical NLP tasks 2024
+SurgeryLLM ong2024surgeryllm GPT4All anand2023gpt4all with Chroma DB Surgical decision support 2024
+Myers et al. Myers2025 LLM2Vec (for comparison) llm2vec Clinical Information retrieval 2024
 Type 5: Domain adaptation retriever
-Self-BioRAG 106 MedCPT 33 Medical QA 2024
-WeiseEule 101 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
-CliniqIR 52 MedCPT and BM25 Diagnostic decision support 2024
-MEDRAG 101 BM25, SPECTER 22, Contriever 57and MedCPT Biomedical RAG Tool 2024
-RAMIE 107 MedCPT, Contriever and BMRetriever 55 Biomedical IR about dietary supplements 2025
+Self-BioRAG jeong2024improving MedCPT jin2023medcpt Medical QA 2024
+WeiseEule Aftab2024 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
+CliniqIR Abdullahi2024 MedCPT and BM25 Diagnostic decision support 2024
+MEDRAG Aftab2024 BM25, SPECTER specter2020cohan, Contriever izacard2021contrieverand MedCPT Biomedical RAG Tool 2024
+RAMIE zhan2025ramie MedCPT, Contriever and BMRetriever xu2024bmretriever Biomedical IR about dietary supplements 2025
 Type 2: Commercial embedding APIs.
@@
 As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities.
-While OpenAI’s offerings dominate, alternatives are emerging.
-The RAP system 92 utilizes Amazon’s Titan Text Embeddings v2 93 for nutrition-related question answering.
-Other commercial APIs such as Anthropic’s Voyage-2 108 and Google’s Vertex AI embeddings 109, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
+While OpenAI’s offerings currently dominate, alternative solutions are also emerging.
+For example, the RAP system azimi2025evaluation utilizes Amazon’s Titan Text Embeddings v2 aws_titan_embedding for nutrition-related question answering, and other commercial APIs, such as Anthropic’s Voyage-2 anthropic_embeddings and Google’s Vertex AI embeddings google_vertexai_embeddings, though not represented in Table LABEL:tab:Classification, can also be employed in biomedical RAG applications
 Type 3: Domain adaptation PLMs.
-Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 96, 98.
-These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 110.
-By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 110.
-As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 98, BioBERT 96 and BioGPT 102, which have been widely adopted in biomedical RAG systems.
-For example, PM-Search 24 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision.
-Similarly, PubMedBERT has been successfully deployed in LADER 97 for biomedical literature retrieval and in KG-RAG 100 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks.
-More recent systems like WeiseEule 101 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 33 to achieve enhanced performance in biomedical QA.
+Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 10.1093/bioinformatics/btz682, pubmedbert.
+These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature Lyu2024LanguageMA.
+By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpora, these models develop a more nuanced understanding of biomedical language, resulting in competitively improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts Lyu2024LanguageMA.
+As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT pubmedbert, BioBERT 10.1093/bioinformatics/btz682, and BioGPT luo2022biogpt, which have been widely adopted in biomedical RAG systems.
+For instance, PM-Search Jin2022 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision.
+Similarly, PubMedBERT has been successfully deployed in LADER LADER for biomedical literature retrieval and in KG-RAG Soman2024 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks.
+More recent systems like WeiseEule Aftab2024 compare multiple biomedical models in their study, including BioBERT, BioGPT, and MedCPT jin2023medcpt, to achieve enhanced performance in biomedical QA.
 Type 4: LLM-based embedding.
 Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context.
-With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 12.
-Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches.
-As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 4, which utilizes MedLLaMA 13b 111, SurgeryLLM 28, which employs GPT4All 103, and the comparative study by Myers et al.
-104, which evaluates LLM2Vec 105 for clinical information retrieval applications.
+With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text Zhao2023LLMSurvey.
+Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings.
+As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG li2025biomedrag, which utilizes MedLLaMA-13B wu2023pmc; SurgeryLLM ong2024surgeryllm, which employs GPT4All anand2023gpt4all; and the comparative study by Myers et al.
+Myers2025, which evaluates LLM2Vec llm2vec for clinical information retrieval applications.
 However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations.
-Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 112, Landmark Embedding 113, which provide promising dense retrieval alternatives in biomedical RAG.
+Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec li-etal-2024-llama2vec, Landmark Embedding luo-etal-2024-landmark, which provide promising dense retrieval alternatives in biomedical RAG.
 Type 5: Domain adaptation retrievers.
@@
 Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed.
-As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 57, MedCPT 33, SPECTER 22, BMRetriever 55, and the MEDRAG toolkit 4.
-These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 106, 52, 107.
+As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever izacard2021contriever, MedCPT jin2023medcpt, SPECTER specter2020cohan, BMRetriever xu2024bmretriever, and the MEDRAG toolkit li2025biomedrag.
+These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies jeong2024improving, Abdullahi2024, zhan2025ramie.
 For example, Jeong et al.
-106 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach.
-Similarly, the MEDRAG framework 101 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool.
-RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 107.
+jeong2024improving proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach.
+Similarly, the MEDRAG framework Aftab2024 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool.
+RAMIEzhan2025ramie utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments zhan2025ramie.
 These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
-Dense Retriever Infrastructure.
-Vector stores provide an infrastructure element for dense retrievers in semantic search systems 114.
-These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers.
-Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems.
-The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches.
-Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3.
-In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
 Table 4: Common vector stores in biomedical RAG
 Store Description
-Chroma 115 AI-native open-source vector database with built-in functionality
-Faiss 116 Library for efficient similarity search and clustering of dense vectors
-Pinecone 117 Production-ready vector database for similarity search at scale
-Weaviate 118 Graph-based vector search engine with semantic capabilities
-HNSW 78 Hierarchical navigable small world graphs for approximate nearest neighbor search
+Chroma chroma2023 AI-native open-source vector database with built-in functionality
+Faiss faiss2023 Library for efficient similarity search and clustering of dense vectors
+Pinecone pinecone_overview Production-ready vector database for similarity search at scale
+Weaviate weaviate_developers Graph-based vector search engine with semantic capabilities
+HNSW malkov2016efficient Hierarchical navigable small world graphs for approximate nearest neighbor search
+Dense Retriever Infrastructure.
+Vector stores provide an infrastructure element for dense retrievers in semantic search systems han2023comprehensive.
+These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers.
+Vector databases enable rapid retrieval of semantically similar documents at scale, thereby improving the efficiency and reducing the response time of biomedical RAG systems.
+The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches.
+Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we compiled the most commonly used vector databases in Table 4, among which FAISS and Chroma emerge as the most popular choices, being utilized 4 and 3 times, respectively.
 Knowledge Graph Enhanced Retrieval
-Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods.
-Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability119.
-Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source.
-As we can notice, the application areas vary significantly, including medical question answering 100, knowledge extraction 120, and diagnosis prediction 17.
+Knowledge graph retrievers provide structured, interlinked domain knowledge in biomedical RAG, helping to address several limitations of traditional retrieval methods.
+Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainabilitypeng2024graph.
+Table 5 presents recent biomedical RAG studies that incorporate knowledge graphs, outlining their application domains and the knowledge graph sources they employ.
+The application areas covered by these systems include medical question answering Soman2024, knowledge extraction DiMaria2024, and diagnosis prediction Gao2025.
 Table 5: Recent biomedical RAG studies based on knowledge graphs
 Method Task Knowledge Graph Used Year
-KG-RAG 100 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
-HEALIE 29 Personalized medical content generation HEALIE KG (customized) 2024
-KRAGEN 121 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
-Ascle 30 Various medical text generation tasks UMLS 2024
-DALK 122 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
-Gilbert et al. 123 Medical information structuring and interlinking Not specified 2024
-NetMe 2.0 120 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
-NEKO 124 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
-DR.KNOWS 17 Diagnosis prediction from EHRs UMLS 2025
-ESCARGOT 125 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
-As shown in Fig 2, the integration of knowledge graphs with LLMs in biomedical RAG applications can be analyzed through four key components:
-Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
+KG-RAG Soman2024 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
+HEALIE Kakalou2024 Personalized medical content generation HEALIE KG (customized) 2024
+KRAGEN Matsumoto2024 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
+Ascle Ascle Various medical text generation tasks UMLS 2024
+DALK li2024dalk Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
+Gilbert et al. Gilbert2024 Medical information structuring and interlinking Not specified 2024
+NetMe 2.0 DiMaria2024 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
+NEKO XIAO202560 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
+DR.KNOWS Gao2025 Diagnosis prediction from EHRs UMLS 2025
+ESCARGOT Matsumoto2025 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
+Figure 3: The Framework of Knowledge Graphs-Based Biomedical RAG
+To further clarify how knowledge graphs are incorporated into biomedical RAG systems, Fig. 3 illustrates the integration of knowledge graphs with LLMs through four key components:
 Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories.
-NetMe 2.0 120 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 124 mines PubMed literature for synthetic biology knowledge graphs.
-Differently, KG-RAG 100 utilizes the pre-existing SPOKE biomedical knowledge graph.
+For example, NetMe 2.0 DiMaria2024 constructs biomedical knowledge graphs on-the-fly using OntoTagMe for entity linking, while NEKO XIAO202560 mines PubMed literature to build knowledge graphs for synthetic biology.
+By contrast, KG-RAG Soman2024 utilizes the pre-existing SPOKE biomedical knowledge graph.
 Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph.
-ESCARGOT 125 employs Cypher queries for precise retrieval, KRAGEN 121 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 17 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
+For instance, ESCARGOT Matsumoto2025 employs Cypher queries for precise retrieval, KRAGEN Matsumoto2024 vectorizes knowledge graphs for semantic searching, and DR.KNOWS Gao2025 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
 Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query.
-DR.KNOWS 17 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 121 applies graph-of-thoughts for multi-hop relationships, and HEALIE 29 extracts paths incorporating both medical knowledge and socio-economic factors.
-  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency.
-KG-RAG 100 implements token-optimized retrieval to reduce consumption, ESCARGOT 125 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 120 transforms graph data into natural language prompts via its Graph-RAG module.
+Specifically, DR.KNOWS Gao2025 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN Matsumoto2024 applies graph-of-thoughts for multi-hop relationships, and HEALIE Kakalou2024 extracts paths incorporating both medical knowledge and socio-economic factors.
+Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts in ways that maximize factual accuracy while optimizing token usage and computational efficiency.
+For example, KG-RAG Soman2024 implements token-optimized retrieval to reduce consumption, whereas ESCARGOT Matsumoto2025 selectively integrates knowledge through a dynamic Graph of Thoughts.
+In addition, NetMe 2.0 DiMaria2024 transforms graph data into natural language prompts via its Graph-RAG module.
 Multi-modal Retrieval
 Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements.
-These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 126.
-As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 32, prescription label interpretation 127, lung cancer staging 128, digital pathology analysis 129, and automated radiology report generation 130, 31, 131, 132, 133.
+These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches abootorabi2025askmodalitycomprehensivesurvey.
+As shown in Table 6, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection Upadhyaya2024, prescription label interpretation Thetbanthad2025, lung cancer staging Tozuka2024, digital pathology analysis Omar2024, and automated radiology report generation pmlr-v219-ranjit23a, Sun2024FactAware, xia2024mmedrag, Xia2024RULE, Yang2025STREAM.
 Figure 4: Architectural workflow of multimodal biomedical RAG system
@@
 Model Task Modalities Year
-Ranjit et al. 130 Chest X-Ray report generation Chest X-ray and text 2023
-FactMM-RAG 31 Radiology report generation X-Ray and text 2024
-MMed-RAG 131 Medical VQA & Report generation Medical images and text 2024
-MEAG 32 Amblyopia diagnosis Eye tracking and text 2024
-Tozuka et al.  128 Lung cancer staging CT findings and TNM classifications 2024
-GPT4DFCI-RAG 129 Digital pathology QA Pathology images and text 2024
-Raminedi et al. 19 Radiology report generation X-ray images and text 2024
-RULE 132 Medical VQA & Report generation Medical images and text 2024
-Thetbanthad et al. 127 Prescription label identification Image (labels) and text 2025
-Hu et al. 134 Pathology report generation Whole slide images and text 2025
-STREAM 133 Chest X-ray report generation Chest X-ray and text 2025
+Ranjit et al. pmlr-v219-ranjit23a Chest X-Ray report generation Chest X-ray and text 2023
+FactMM-RAG Sun2024FactAware Radiology report generation X-Ray and text 2024
+MMed-RAG xia2024mmedrag Medical VQA & Report generation Medical images and text 2024
+MEAG Upadhyaya2024 Amblyopia diagnosis Eye tracking and text 2024
+Tozuka et al. Tozuka2024 Lung cancer staging CT findings and TNM classifications 2024
+GPT4DFCI-RAG Omar2024 Digital pathology QA Pathology images and text 2024
+Raminedi et al. Raminedi2024MultimodalTransformer Radiology report generation X-ray images and text 2024
+RULE Xia2024RULE Medical VQA & Report generation Medical images and text 2024
+Thetbanthad et al. Thetbanthad2025 Prescription label identification Image (labels) and text 2025
+Hu et al. Hu2025 Pathology report generation Whole slide images and text 2025
+STREAM Yang2025STREAM Chest X-ray report generation Chest X-ray and text 2025
 Figure 4 illustrates the architectural workflow of a multimodal biomedical RAG system.
 The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture.
-While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 126.
-The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 135 for images, specialized encoders for time-series data 135); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge.
-For instance, Ranjit et al.130, FactMM-RAG 31 and STREAM 133 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report.
-MMed-RAG 131 and RULE 132 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation.
-The MEAG 32 system uses specialized encoders for eye movement data alongside text, Raminedi et al.
-129 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al.
-127 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
+While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs abootorabi2025askmodalitycomprehensivesurvey.
+The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers chen2021outperform for images, specialized encoders for time-series data chen2021outperform); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge.
+For instance, Ranjit et al.pmlr-v219-ranjit23a, FactMM-RAG Sun2024FactAware and STREAM Yang2025STREAM use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report.
+MMed-RAG xia2024mmedrag and RULE Xia2024RULE build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation.
+The MEAG Upadhyaya2024 system uses specialized encoders for eye movement data alongside text, Raminedi et al.
+Omar2024 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al.
+Thetbanthad2025 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
 Reranker
-Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 45.
-Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component.
+Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models sivarajkumar2024clinical.
+Positioned between retrievers and generators, rerankers apply appropriate ranking algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component.
 By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
 Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies.
-Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 52 and MEDRAG 101 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 83 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity.
-Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 101 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval.
-Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 34 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 136 utilizes a specialized cross-encoder model known as BGE M3-Embedding 137 to enhance clinical document relevance scoring.
-Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 33 trains a specific reranker denoted as CrossEnc.
-Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 77 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 138 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7.
+Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR Abdullahi2024 and MEDRAG Aftab2024 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI toro2024dragonai employs Maximal Marginal Relevance (MMR) to balance relevance with diversity.
+In addition, WeiseEule Aftab2024 introduces a keyword frequency–based ranking system tailored for biomedical literature retrieval.
+Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT jia2024podgpt uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system adejumo2024retrieval utilizes a specialized cross-encoder model known as BGE M3-Embedding chen2024m3embedding to enhance clinical document relevance scoring.
+Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT jin2023medcpt trains a specific reranker denoted as CrossEnc.
+Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA darnell2024biomedical employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC craig2024lmrac implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7.
 (on a 10 point scale).
 Finally, hybrid approaches combine multiple strategies to enhance performance.
-For example, Clinfo.ai 14 combines LLM classification with BM25 to re-rank retrieved articles.
+For example, Clinfo.ai Lozano2024ClinfoAI combines LLM classification with BM25 to re-rank retrieved articles.
 Generation
 The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries.
-In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 139.
+In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions lee2024enhancing.
 This section examines various generation models employed in recent biomedical RAG studies.
@@
 As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception.
-This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 12.
-Among open-source LLMs, LLaMa 2 140 is the most popular, with 15 biomedical applications built upon it.
-LLaMa 3 38, Mistral 7B 141, and T5 142 are also widely used.
+This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency Zhao2023LLMSurvey.
+Among open-source LLMs, LLaMa 2 HugoTouvron2023LLaMA2 is the most popular, with 15 biomedical applications built upon it.
+LLaMa 3 MetaAI2024LLaMA3, Mistral 7B mistral7b_2024, and T5 2020t5 are also widely used.
 These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems.
-Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 40 and Gemma 3 143.
+Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 mistral_small_24b_2024 and Gemma 3 gemma-3-27b-it.
 Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
 For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance.
-Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 85, 144, 15, thereby offering viable alternatives to ChatGPT.
+Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications wang2024diabetesrag, hewitt2024llmneuropathology, Lammert2024ExpertLLM, thereby offering viable alternatives to ChatGPT.
 Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
 In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications.
-Recent studies 145, 107, 146 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance.
+Recent studies ozaki2024confidence, zhan2025ramie, alkaeed2025open have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance.
 Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
@@
 Open-Source LLMs
-T5 142 0.06-11B Encoder-Decoder VAIV 53, CLEAR 99, DR.KNOWS 17
-ChatGLM3 147 6B Decoder-only COPD 148, ChatZOC 149
-LLaMA 2 140 7-70B Decoder-only Guo et al.
-27, KREIMEYER et al.
-150, CaLM 26, MMRAG 151, Yu et al.
-152, RAMIE 107, EyeGPT 153, KG-RAG 100, JMLR 3, BiomedRAG 4, Self-BioRAG 106, CaLM 26, Du et al.
-154, SurgeryLLM 28, Kreimeyer et al.
-150
-LLaMA 3 38 8-70B Decoder-only MMRAG 151, Fatharanihttps et al. 155, RAMIE 107, Woo et al. 156, RAG-HPO 72, i-MedRAG 157, Hewitt et al. 144, PodGPT 34, SurgeryLLM 28
-Phi-3 Mini 158 3.8B Decoder-only Fatharanihttps et al. 155
-Mistral 7B 141 7B Decoder-only Boulos et al. 159, RAMIE 107, Woo et al. 156, RAGPR 80, LITURAt 62, Kreimeyer et al. 150
-QWen 160 32B Decoder-only NEKO 124, Thetbanthad 127
-Baichuan 161 7-13B Decoder-only ChatZOC 149
-Falcon 162 7-180B Decoder-only AskFDALabel 42, CaLM 26
-Zephyr 163 7B Decoder-only Moser et al. 164
-AceGPT 165 7-13B Decoder-only InfectA-Chat 94
-Gemma 166 1-27B Decoder-only PodGPT 34
-Deepseek R1 41 671B Decoder-only Feng et al. 167
+T5 2020t5 0.06-11B Encoder-Decoder VAIV Kim2024, CLEAR lopez2025clinical, DR.KNOWS Gao2025
+ChatGLM3 glm2024chatglm 6B Decoder-only COPD Copd-ChatGLM, ChatZOC luo2024retrieval
+LLaMA 2 HugoTouvron2023LLaMA2 7-70B Decoder-only Guo et al.
+Guo2024, KREIMEYER et al.
+kreimeyer2024rag, CaLM parmanto2024calm, MMRAG zhan2025mmrag, Yu et al.
+pmlr-v225-yu23b, RAMIE zhan2025ramie, EyeGPT chen2024eyegpt, KG-RAG Soman2024, JMLR Wang2024JMLR, BiomedRAG li2025biomedrag, Self-BioRAG jeong2024improving, CaLM parmanto2024calm, Du et al.
+du2024cognitive, SurgeryLLM ong2024surgeryllm, Kreimeyer et al.
+kreimeyer2024rag
+LLaMA 3 MetaAI2024LLaMA3 8-70B Decoder-only MMRAG zhan2025mmrag, Fatharanihttps et al.
+fatharani2025pharmacogenomics, RAMIE zhan2025ramie, Woo et al.
+woo2025custom, RAG-HPO garcia2024deepphenotyping, i-MedRAG xiong2025iterative, Hewitt et al.
+hewitt2024llmneuropathology, PodGPT jia2024podgpt, SurgeryLLM ong2024surgeryllm
+Phi-3 Mini phi3mini4k_2024 3.8B Decoder-only Fatharanihttps et al. fatharani2025pharmacogenomics
+Mistral 7B mistral7b_2024 7B Decoder-only Boulos et al. kamel2024chat, RAMIE zhan2025ramie, Woo et al. woo2025custom, RAGPR zheng2025integrating, LITURAt Peasley2025, Kreimeyer et al. kreimeyer2024rag
+QWen qwen32b_2024 32B Decoder-only NEKO XIAO202560, Thetbanthad Thetbanthad2025
+Baichuan wang2025baichuan 7-13B Decoder-only ChatZOC luo2024retrieval
+Falcon almazrouei2023falconseries 7-180B Decoder-only AskFDALabel llm-regulatory-framework, CaLM parmanto2024calm
+Zephyr tunstall2023zephyr 7B Decoder-only Moser et al. moser2024synthetic
+AceGPT huang2024acegpt 7-13B Decoder-only InfectA-Chat selcuk2025infecta
+Gemma team2024gemma 1-27B Decoder-only PodGPT jia2024podgpt
+Deepseek R1 deepseekr1_2024 671B Decoder-only Feng et al. feng2025retrieval
 Commercial LLMs
-ChatGPT-3.5/4/4o 35 Proprietary Decoder-only Woo et al.
-156, RAGPR 80, ChatZOC 149, KG-RAG 100, Clinfo.ai 14, Yu et al.
-152, ChatENT 89, CaLM 26, Rau et al.
-168, Endo-chat 88, BiomedRAG 4, Lu 18, Tan 169, GNQA 77, Puts 170, DRAGON-AI 83, Choi 171, RISE 85, Du et al.
-154,i-MedRAG 157 , Hewitt et al.
-144, RECTIFIER 91, Kainer et al.
-95, Gong et al.
-49, VAIA 53, RareDxGPT 16, Selcuk et al.
-94, FAVOR-GPT 87, Markey et al.
-172,
-Claude-3.5 36 Proprietary Decoder-only Woo et al. 156, RISE 85, Hewitt et al. 144
-Gemini 37 Proprietary Multimodal Upadhyaya et al. 32, Tozuka et al. 128, MEREDITH 15
+ChatGPT-3.5/4/4o OpenAI2025TextGeneration Proprietary Decoder-only Woo et al.
+woo2025custom, RAGPR zheng2025integrating, ChatZOC luo2024retrieval, KG-RAG Soman2024, Clinfo.ai Lozano2024ClinfoAI, Yu et al.
+pmlr-v225-yu23b, ChatENT long2024chatent, CaLM parmanto2024calm, Rau et al.
+rau2024retrieval, Endo-chat fu2024application, BiomedRAG li2025biomedrag, Lu Lu2025BoostingGPT, Tan an2025chatgpt, GNQA darnell2024biomedical, Puts puts2025icd10, DRAGON-AI toro2024dragonai, Choi Choi2024, RISE wang2024diabetesrag, Du et al.
+du2024cognitive,i-MedRAG xiong2025iterative , Hewitt et al.
+hewitt2024llmneuropathology, RECTIFIER unlu2024retrieval, Kainer et al.
+kainer2025effectiveness, Gong et al.
+Gong2024, VAIA Kim2024, RareDxGPT Zelin2024RareDisease, Selcuk et al.
+selcuk2025infecta, FAVOR-GPT li2024favorgpt, Markey et al.
+markey2025rags,
+Claude-3.5 Anthropic2025GettingStarted Proprietary Decoder-only Woo et al. woo2025custom, RISE wang2024diabetesrag, Hewitt et al. hewitt2024llmneuropathology
+Gemini GoogleGemini2025 Proprietary Multimodal Upadhyaya et al. Upadhyaya2024, Tozuka et al. Tozuka2024, MEREDITH Lammert2024ExpertLLM
 Biomedical Specialized LLMs
-MedAlpaca 173 7-13B LLaMA-based RAMIE 107
-PMC-LLaMA 174 7B LLaMA-based RAMIE 107, BiomedRAG 4, Ozaki et al. 145
-BioMistral 175 7B Mistral-based RAMIE 107, pRAGe 176
-MEDITRON 177 7-70B LLaMA-based Ozaki et al. 145, Alkaeed et al. 146
-Meerkat 178 7B Mistral-based Sohn et al. 179
+MedAlpaca han2023medalpaca 7-13B LLaMA-based RAMIE zhan2025ramie
+PMC-LLaMA Wu2024PMC-LLaMA 7B LLaMA-based RAMIE zhan2025ramie, BiomedRAG li2025biomedrag, Ozaki et al. ozaki2024confidence
+BioMistral labrak2024biomistral 7B Mistral-based RAMIE zhan2025ramie, pRAGe buhnila2024retrieve
+MEDITRON chen2023meditron70b 7-70B LLaMA-based Ozaki et al. ozaki2024confidence, Alkaeed et al. alkaeed2025open
+Meerkat kim2024small 7B Mistral-based Sohn et al. sohn2024rationale
 Domain Adaptation
-Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 180.
+Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications siriwardhana2023improving.
 This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models.
@@
 Retriever Stage Adaptation.
-Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 57 and MedCPT 33) enable the development of more semantically rich biomedical embeddings.
+Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever izacard2021contriever and MedCPT jin2023medcpt) enable the development of more semantically rich biomedical embeddings.
 These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships.
-Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 55) and domain-specific architectural modifications (exemplified by LADER 97 with PubMedBERT initialization).
+Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER xu2024bmretriever) and domain-specific architectural modifications (exemplified by LADER LADER with PubMedBERT initialization).
 These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
 Reranker Stage Adaptation.
-The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 33) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 4).
-Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage.
-These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria.
-Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
+The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker jin2023medcpt) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer li2025biomedrag).
+Notably, the aforementioned Model-based rerankers and LLM-based rerankers can be conceptualized as domain adaptation strategies operationalized within the reranking stage.
+These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, improve the system’s ability to prioritize documents by capturing subtle domain-specific signals that general models often fail to detect, thereby enhancing retrieval precision for specialized medical queries.
 LLM Stage Adaptation.
-At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 6, SeRTS 56), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 42), and continued pre-training (RALL 27, ChatENT 89) enables more accurate synthesis and explication of biomedical information.
-JLMR 3 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions.
+At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE frisoni2024generate, SeRTS Hu2024), parameter-efficient approaches like Low-Rank Adaptation (LoRA)hu2022lowrank (implemented in AskFDALabel llm-regulatory-framework), and continued pre-training (RALL Guo2024, ChatENT long2024chatent) enables more accurate synthesis and explication of biomedical information.
+JLMR Wang2024JMLR joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions.
 These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
-This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns essential for healthcare and life sciences applications.
-Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, ultimately advancing the state-of-the-art in biomedical information retrieval and synthesis.
-Datasets
+This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns required for healthcare and life sciences applications.
+Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, contributing to ongoing progress in biomedical information retrieval and synthesis.
+Datasets and Evaluation
 The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets.
-In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task datasets.
+In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task benchmarks, along with their corresponding evaluation metrics.
 This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
 Knowledge Source Datasets
-The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 181, 1.
+The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources gao2023retrieval, Fan2024RAG.
 These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content.
-High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 182, 89, 106.
+High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices huang2024toolcalling, long2024chatent, jeong2024improving.
 Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology.
 To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format.
-This integration enables systems to comprehensively address the multifaceted nature of medical queries 101, 55.
+This integration enables systems to comprehensively address the multifaceted nature of medical queries Aftab2024, xu2024bmretriever.
 Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies.
-Among these sources, PubMed, UMLS 183, PMC, and Medical textbooks 184 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections.
-Additionally, we notice that several studies 171, 154, 80 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 50, 185, 76, 92, 165, 186 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
+Among these sources, PubMed, UMLS bodenreider2004umls, PMC, and Medical textbooks jin2020disease demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections.
+Additionally, we notice that several studies Choi2024, du2024cognitive, zheng2025integrating have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies Lee2024, masanneck2025evaluating, xu2024evaluation, azimi2025evaluation, huang2024acegpt, sevgi2024medical focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
 Table 8: Comprehensive Medical Knowledge Sources
 Source Type Description Studies
-MIMIC-IV 187 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 164
-PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 14, MEDRAG 61, Self-BioRAG 106, JMLR 3, Ascle 30, RISE 85, MEREDITH 15, LITURAt 62, DR.KNOWS 17, Aftab 101, CliniqIR 52, PM-Search 24, Kreimeyer 150, Kim 53
-PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 106, Soong et al. 188, PodGPT 34
-TripClick 189 Literature Biomedical literature search and retrieval dataset LADER 97
-GNQA 77 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 77
-UMLS 183 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 4, CLEAR 99, Guo 27, CliniqIR 52
-ICD-10 190 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 170
-HPO 191 Knowledge Base Human phenotype ontology Albayrak et al. 192, RAG-HPO 72, Kainer 95
-MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 55
-AlzKB 125 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 125, KRAGEN 121
-DrugBank 193 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 55, Kim 53
-SPOKE 194 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 100
-COD 149 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 149
-HEALIE KG29 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 29
-ADRD Caregiving 26 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 26
-StatPearls 23 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 61, i-MedRAG 157
-Medical Textbooks 184 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 61, Self-BioRAG 106, JMLR 3, i-MedRAG 157
-Ophthalmology textbooks 153 Educational 14 specialized ophthalmology textbooks EyeGPT 153
-CheXpert 195 Image-Report 224,316 chest radiographs with associated reports CLEAR 99
+MIMIC-IV johnson2023mimic EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. moser2024synthetic
+PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai Lozano2024ClinfoAI, MEDRAG xiong-etal-2024-benchmarking, Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, RISE wang2024diabetesrag, MEREDITH Lammert2024ExpertLLM, LITURAt Peasley2025, DR.KNOWS Gao2025, Aftab Aftab2024, CliniqIR Abdullahi2024, PM-Search Jin2022, Kreimeyer kreimeyer2024rag, Kim Kim2024
+PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG jeong2024improving, Soong et al. soong2024improving, PodGPT jia2024podgpt
+TripClick rekabsaz2021tripclick Literature Biomedical literature search and retrieval dataset LADER LADER
+GNQA darnell2024biomedical Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA darnell2024biomedical
+UMLS bodenreider2004umls Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG li2025biomedrag, CLEAR lopez2025clinical, Guo Guo2024, CliniqIR Abdullahi2024
+ICD-10 cms2023icd10pcs Knowledge Base International statistical classification of diseases and related health problems Puts et al. puts2025icd10
+HPO jsl2024hpo Knowledge Base Human phenotype ontology Albayrak et al. albayrak2024hpo, RAG-HPO garcia2024deepphenotyping, Kainer kainer2025effectiveness
+MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER xu2024bmretriever
+AlzKB Matsumoto2025 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT Matsumoto2025, KRAGEN Matsumoto2024
+DrugBank Wishart2006 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER xu2024bmretriever, Kim Kim2024
+SPOKE morris2023spoke Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG Soman2024
+COD luo2024retrieval Knowledge Base Comprehensive ophthalmic dataset ChatZOC luo2024retrieval
+HEALIE KGKakalou2024 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. Kakalou2024
+ADRD Caregiving parmanto2024calm Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM parmanto2024calm
+StatPearls Bashir2024 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG xiong-etal-2024-benchmarking, i-MedRAG xiong2025iterative
+Medical Textbooks jin2020disease Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG xiong-etal-2024-benchmarking, Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, i-MedRAG xiong2025iterative
+Ophthalmology textbooks chen2024eyegpt Educational 14 specialized ophthalmology textbooks EyeGPT chen2024eyegpt
+CheXpert irvin2019chexpert Image-Report 224,316 chest radiographs with associated reports CLEAR lopez2025clinical
 Medical Task Datasets
-Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 2, 10.
+Medical task datasets represent important benchmarking instruments for evaluating RAG systems within healthcare applications he2023survey, liu2024survey.
 These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
@@
 Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems.
-Apart from works that make use of public biomedical datasets, several research groups 81, 90, 85 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
+Apart from works that make use of public biomedical datasets, several research groups Fukushima2025, steybe2025evaluation, wang2024diabetesrag in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
 Table 9: Comprehensive Biomedical Tasks Datasets
 Dataset Year Task Studies
-ChemProt 196 2010 Information Extraction BiomedRAG 4, ChemProt
-DDI 197 2013 Information Extraction BiomedRAG 4, ChemProt
-NCBI 198 2014 Entity Recognition GeneGPT 199, NetMe 120, RT 74
-BioASQ 200 2015 Information Retrieval BMRETRIEVER 55, SeRTS 56
-MIMIC-III 20 2016 Text Summarization DR.KNOWS 17, CliniqIR 52
-BC5CDR 201 2016 Entity Recognition RT 74
-DC3 202 2019 Diagnostic classification CliniqIR 52
-MIMIC-CXR 203 2019 Text Summarization Ascle 30
-MedQuAD 204 2019 QA Ascle 30
-PubMedQA 205 2019 Multiple-choice BMRETRIEVER 55, Liévin et al. 206, PodGPT 34
-MMLU 207 2021 Multiple-choice Self-BioRAG 106, JMLR 3, i-MedRAG 157, InfectA-Chat 94, PodGPT 34
-MedQA 184 2021 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, i-MedRAG 157, PodGPT 34
-HealthSearchQA 208 2022 QA
-MedMCQA 209 2022 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, EyeGPT 153, PodGPT 34
-ChatDoctor 210 2023 Dialogue EyeGPT 153
-MedAlpaca 173 2023 Dialogue EyeGPT 153
-GeneTuring 211 2023 Genomics QA GeneGPT 199
-GIT 212 2023 Information Extraction BiomedRAG 4
-MIRAGE 61 2024 Multiple-choice & QA MEDRAG 61
-MedExpQA 51 2024 Multilingual Medical QA MedExpQA 51, PodGPT 34
-PubMedRS-200 14 2024 Medical QA Clinifo.ai 14
-MedicineQA 182 2024 Multi-round dialogue RagPULSE 182
-CELLS 27 2024 Lay language generation Guo 27
-Applications
+ChemProt taboureau2010chemprot 2010 Information Extraction BiomedRAG li2025biomedrag, ChemProt
+DDI segura2013semeval 2013 Information Extraction BiomedRAG li2025biomedrag, ChemProt
+NCBI dogan2014ncbi 2014 Entity Recognition GeneGPT jin2024genegpt, NetMe DiMaria2024, RT li2024rt
+BioASQ Tsatsaronis2015 2015 Information Retrieval BMRETRIEVER xu2024bmretriever, SeRTS Hu2024
+MIMIC-III johnson2016mimic 2016 Text Summarization DR.KNOWS Gao2025, CliniqIR Abdullahi2024
+BC5CDR li2016biocreative 2016 Entity Recognition RT li2024rt
+DC3 eickhoff2019dc3 2019 Diagnostic classification CliniqIR Abdullahi2024
+MIMIC-CXR johnson2019mimic 2019 Text Summarization Ascle Ascle
+MedQuAD BenAbacha-BMC-2019 2019 QA Ascle Ascle
+PubMedQA jin2019pubmedqa 2019 Multiple-choice BMRETRIEVER xu2024bmretriever, Liévin et al. Lievin2024, PodGPT jia2024podgpt
+MMLU hendryckstest2021 2021 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, i-MedRAG xiong2025iterative, InfectA-Chat selcuk2025infecta, PodGPT jia2024podgpt
+MedQA jin2020disease 2021 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, Liévin et al. Lievin2024, i-MedRAG xiong2025iterative, PodGPT jia2024podgpt
+HealthSearchQA singhal2022large 2022 QA
+MedMCQA pal2022medmcqa 2022 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, Liévin et al. Lievin2024, EyeGPT chen2024eyegpt, PodGPT jia2024podgpt
+ChatDoctor li2023chatdoctor 2023 Dialogue EyeGPT chen2024eyegpt
+MedAlpaca han2023medalpaca 2023 Dialogue EyeGPT chen2024eyegpt
+GeneTuring hou2023geneturing 2023 Genomics QA GeneGPT jin2024genegpt
+GIT li2023petailor 2023 Information Extraction BiomedRAG li2025biomedrag
+MIRAGE xiong-etal-2024-benchmarking 2024 Multiple-choice & QA MEDRAG xiong-etal-2024-benchmarking
+MedExpQA ALONSO2024102938 2024 Multilingual Medical QA MedExpQA ALONSO2024102938, PodGPT jia2024podgpt
+PubMedRS-200 Lozano2024ClinfoAI 2024 Medical QA Clinifo.ai Lozano2024ClinfoAI
+MedicineQA huang2024toolcalling 2024 Multi-round dialogue RagPULSE huang2024toolcalling
+CELLS Guo2024 2024 Lay language generation Guo Guo2024
+Evaluation Metrics
+The biomedical RAG systems reviewed in this survey typically follow a three-stage pipeline, consisting of (1) retrieval, (2) reranking, and (3) generation.
+Consequently, the evaluation methodologies used across the surveyed studies can also be naturally organized according to these stages.
+Because each stage involves a distinct functional component of the RAG architecture, the choice of evaluation metrics naturally varies across stages.
+Moreover, as shown in Table LABEL:tab:biomedical-comprehensive, biomedical RAG systems are applied to a wide range of downstream tasks—such as NER, medical QA, dialogue, and clinical summarization—and the selection of evaluation metrics further depends on the specific characteristics of these tasks.
+In what follows, we review the evaluation metrics used in biomedical RAG research by following this pipeline-based structure.
+Metrics for Retrieval
+The retrieval stage evaluates how effectively a system surfaces and orders clinically relevant documents from large-scale biomedical corpora, with emphasis on ranking quality, relevance precision, and coverage of essential evidence.
+More broadly, evaluation in this stage follows an information-retrieval paradigm, where the goal is to quantify how well the retrieved set and its ordering align with the ground-truth relevance distribution.
+We review the core retrieval metrics commonly adopted in biomedical RAG systems—namely, Precision@k, Recall@k, Mean Average Precision (MAP@k), and normalized Discounted Cumulative Gain (nDCG@k)—and provide their formal definitions.
+  1. (1)
+    Precision@k
+    $\text{Precision@k}=\frac{\sum_{i=1}^{k}\mathbf{1}[\text{doc}_{i}\text{ is relevant}]}{k}$ (7)
+  2. (2)
+    Recall@k
+    $\text{Recall@k}=\frac{\sum_{i=1}^{k}\mathbf{1}[\text{doc}_{i}\text{ is relevant}]}{\lvert\text{Relevant Documents}\rvert}$ (8)
+  3. (3)
+    Mean Average Precision (MAP@k)
+    $\text{AP@k}=\frac{1}{R}\sum_{i=1}^{k}\text{Precision@i}\cdot\mathbf{1}[\text{doc}_{i}\text{ is relevant}]$ (9)
+    $\text{MAP@k}=\frac{1}{N}\sum_{j=1}^{N}\text{AP@k}(j)$ (10)
+  4. (4)
+    Normalized Discounted Cumulative Gain (nDCG@k)
+    $\text{DCG@k}=\sum_{i=1}^{k}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$ (11)
+    $\text{nDCG@k}=\frac{\text{DCG@k}}{\text{IDCG@k}}$ (12)
+Metrics for Reranking
+The reranking stage refines the ordering of retrieved documents using more expressive scoring mechanisms (e.g., cross-encoder relevance models).
+Because its objective remains aligned with retrieval—prioritizing the most clinically relevant evidence—biomedical RAG systems typically evaluate reranking with the same IR metrics: Precision@k, Recall@k, MAP@k, and nDCG@k.Notably, in RAG applications, reranking quality is particularly significant because even small improvements in the top-ranked documents can substantially enhance the factual grounding and reliability of the final generated output.
+Metrics for Generation
+The generation stage is evaluated primarily through task-specific criteria, as the correctness and quality of the final output depend heavily on the nature of the downstream biomedical task.
+Because different applications—such as medical QA, clinical summarization, diagnostic classification, and dialogue—impose distinct requirements, the choice of evaluation metrics must be aligned with the characteristics of each task.
+In what follows, we group representative evaluation metrics into four categories: (1) generation metrics; (2) classification metrics; (3) named entity recognition metrics for biomedical IE tasks; and (4) dialogue metrics for assessing multi-turn clinical conversations.
+Generation Metrics assess the quality of free-form textual outputs by measuring aspects such as content correctness, semantic alignment, and linguistic coherence.
+They are applicable to downstream biomedical generation tasks that require producing narrative responses, including medical QA, clinical report generation, and clinical summarization.
+  1. (1)
+    ROUGE-N
+    $\text{ROUGE-N}=\frac{\text{Overlap of N-grams}}{\text{Total N-grams in reference}}$ (13)
+    ROUGE-N measures the n-gram overlap between generated text and the reference text, commonly used for summarization and long-form generation.
+  2. (2)
+    ROUGE-L
+    $\text{ROUGE-L}=\frac{\text{LCS}(X,Y)}{\text{Length of reference}}$ (14)
+    ROUGE-L is based on the Longest Common Subsequence (LCS), capturing sentence-level structural similarity beyond exact n-gram matching.
+  3. (3)
+    BLEU
+    $\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{4}w_{n}\log p_{n}\right)$ (15)
+    BLEU evaluates the precision of n-grams and is widely adopted for machine translation and other structured text generation tasks.
+  4. (4)
+    BERTScore
+    BERTScorebert-score evaluates semantic similarity between a generated text $X$ and a reference text $Y$ by aligning their token embeddings from a pretrained language model.
+Let $e_{x}$ and $e_{y}$ denote contextual embeddings of tokens $x\in X$ and $y\in Y$, respectively.
+The precision $P$ and recall $R$ are defined as
+    $P=\frac{1}{|X|}\sum_{x\in X}\max_{y\in Y}\cos\bigl(e_{x},e_{y}\bigr),\quad R=\frac{1}{|Y|}\sum_{y\in Y}\max_{x\in X}\cos\bigl(e_{y},e_{x}\bigr),$ (16)
+    and the final BERTScore is computed as the $F_{1}$ score:
+    $\text{BERTScore}_{F_{1}}=\frac{2\cdot P\cdot R}{P+R}.$ (17)
+    Compared to lexical overlap metrics, BERTScore offers a more contextualized evaluation by operating in the embedding space.
+  5. (5)
+    BLEURT
+    BLEURTsellam-etal-2020-bleurt is a learned evaluation metric based on pretrained language models fine-tuned on human-rated text pairs.
+Given a generated text $X$ and reference text $Y$, BLEURT encodes both sequences with a transformer encoder and predicts a scalar quality score:
+    $\text{BLEURT}(X,Y)=f_{\theta}\bigl(\text{Enc}(X),\,\text{Enc}(Y)\bigr),$ (18)
+    where $\text{Enc}(\cdot)$ denotes contextual embeddings and $f_{\theta}$ is a regression head trained to approximate human judgments.
+BLEURT captures semantic adequacy and fluency beyond surface-level overlap.
+  6. (6)
+    RaTEScore
+    RaTEScorezhao2024ratescore is an entity-based metric for radiology report generation.
+It extracts medical entities and their relations from both generated and reference reports, matches them with synonym awareness, weights them by clinical importance, and computes an F1-style score.
+This makes the metric focus on factual, clinically meaningful consistency rather than surface wording, achieving higher alignment with expert judgment.
+Classification metrics evaluate the correctness of discrete predictions made by a model, such as assigning labels or choosing among predefined options.
+These metrics are well-suited for tasks like diagnosis classification and multiple-choice question answering.
+  1. (1)
+    Accuracy (ACC) measures the proportion of correctly predicted instances:
+    $\text{ACC}=\frac{\text{\# correct predictions}}{\text{\# total instances}}$ (19)
+  2. (2)
+    Precision, Recall, and F1 for a single class are defined as:
+    $\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN},\quad\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (20)
+  3. (3)
+    Macro-F1 is commonly used in multi-class and multi-label classification settings.
+It computes the F1 score independently for each class and then averages them, giving equal weight to both frequent and rare classes:
+    $\text{Macro-F1}=\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Precision}_{c}\cdot\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}}$ (21)
+  4. (4)
+    Micro-F1 is another metric for multi-class and multi-label classification tasks.
+It aggregates true positives, false positives, and false negatives across all classes before computing a global F1 score, thus favoring frequent classes:
+    $\text{Micro-F1}=\frac{2\cdot\sum_{c=1}^{C}TP_{c}}{2\cdot\sum_{c=1}^{C}TP_{c}+\sum_{c=1}^{C}FP_{c}+\sum_{c=1}^{C}FN_{c}}$ (22)
+AUROC and AUPRC are probability-based evaluation metrics commonly used to assess the performance of binary classifiers across different decision thresholds.
+We notice that they are rarely reported in the surveyed works, as LLMs typically generate free-text outputs or are prompted to select a single label from a predefined set rather than produce explicit class probabilities.
+Named Entity Recognition (NER) metrics measures how accurately a system identifies and labels biomedical entities in text, and are widely used for assessing biomedical information extraction tasks.
+  1. (1)
+    Token-level F1 evaluates NER performance at the token granularity by computing the F1 score over token-wise precision and recall:
+    $\text{Token-F1}=\frac{2\cdot P\cdot R}{P+R}$ (23)
+  2. (2)
+    Entity-level F1 assesses how accurately complete entity spans are identified, considering correct entities, boundary errors, and missed entities:
+    $\text{Entity-F1}=\frac{2\cdot\lvert\text{Correct Entities}\rvert}{2\cdot\lvert\text{Correct Entities}\rvert+\lvert\text{Wrong Boundaries}\rvert+\lvert\text{Missed Entities}\rvert}$ (24)
+Medical Dialogue evaluation is inherently complex, as it must account for both the linguistic quality of generated responses and the effectiveness of completing dialogue-level clinical tasks.
+In practice, evaluation combines automatic metrics with human or LLM-based judgments, since purely automated measures are often insufficient for safety-sensitive biomedical scenarios.
+  1. (1)
+    Generation quality (automatic): Metrics such as ROUGE, BLEU, BERTScore, and DialogRPTgao-etal-2020-dialogue are used to automatically assess lexical overlap, semantic similarity, and response coherence.
+These metrics provide scalable proxies for linguistic quality but do not capture clinical correctness.
+  2. (2)
+    Task success (primarily automatic): Indicators such as task completion rateli2021how and slot accuracyaksu2024granular measure whether the system achieves predefined clinical or informational goals.
+These metrics can be computed automatically when gold annotations are available, although human assessment is sometimes required for semantically defined success criteria.
+  3. (3)
+    Safety and factuality (human/LLM-based): Evaluation of hallucination, factual consistency, and clinical appropriateness typically relies on expert reviewers or medically aligned LLM judges Croxford2025, metric-survey, as automated approaches have limited ability to detect subtle or domain-specific medical errors.
+Automatic hallucination or consistency scores may be used as auxiliary signals, but do not replace human or expert-guided evaluation.
+Together, these complementary evaluation strategies constitute a diverse assessment framework, and surveyed studies typically adopt one or more of these metrics to evaluate the quality of RAG-based medical dialogue systems.
+Applications in biomedicine
 The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval.
-This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 213.
-In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
+This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome liu2025improving.
+In this section, we analyze 5 key domains in which we identified that biomedical RAG systems have demonstrated substantial utility: clinical decision support, medical documentation, precision medicine, education, and clinical research.
 Clinical Decision Support Systems
 Medical Question Answering
-Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 100.
-This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
-Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval.
-Clinfo.ai 14 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence.
-RALL 27 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information.
-Self-BioRAG 106 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 101 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties.
+Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need Soman2024.
+This application is particularly valuable because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
+Recent advancements have produced several novel medical QA systems with distinctive approaches to knowledge retrieval.
+Clinfo.ai Lozano2024ClinfoAI provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence.
+RALL Guo2024 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the growing need for accessible medical information.
+Self-BioRAG jeong2024improving implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule Aftab2024 evaluates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties.
 These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
 Diagnostic and Treatment Decision Support
-Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 214.
+Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis Sutton2020CDSS.
 They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
 Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties.
-Expert-Guided LLMs 15 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data.
-DR.KNOWS 17 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments.
-CliniqIR 52 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 28 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data.
-RECTIFIER 91 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data.
-By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
+Expert-Guided LLMs Lammert2024ExpertLLM support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data.
+DR.KNOWS Gao2025 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments.
+CliniqIR Abdullahi2024 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM ong2024surgeryllm delivers surgical decision support by retrieving relevant procedures, techniques, and complication data.
+RECTIFIER unlu2024retrieval employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data.
+By connecting patient-specific data with evidence-based medical knowledge, these RAG systems help clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
 Rare Disease Identification and Management
-Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 215.
+Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns MartinezDeMiguel2022RareDis.
 RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
@@
 Zelin et al.
-16 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases.
-RAG-HPO 72 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping.
+Zelin2024RareDisease developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases.
+RAG-HPO garcia2024deepphenotyping utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping.
 Yang et al.
-216 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases.
-Some authors 217 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis.
-These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
+Yang2024RDguru created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases.
+A recent study kim2024phenotype has also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis.
+These applications demonstrate how RAG systems can help bridge knowledge gaps in rare disease management by connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
 Clinical Report Generation
-Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 218.
-A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 219.
-This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 220.
-RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 7.
+Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care sloan2024automated.
+A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation hamamci2024ct2rep.
+This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption alfarghaly2021automated.
+RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs xiong2024benchmarking.
 Recent works have demonstrated the effectiveness of RAG in this domain.
-For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 221, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent.
-To enhance interpretability, some authors 222 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability.
-RAG pipelines 223 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors.
-Differently, LaB-RAG 224 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM.
-Lastly, some authors 225 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
+For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 ranjit2023retrieval, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent.
+To enhance interpretability, some authors alam2024towards proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability.
+RAG pipelines bernardi2024report have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors.
+Differently, LaB-RAG song2024lab employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM.
+Lastly, some authors sun2024fact have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
 Precision Medicine Applications
@@
 Genomic medicine involves analyzing genetic data to inform clinical decision-making.
-The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 226.
+The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity burton2012genomic.
 Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval.
-FAVOR-GPT 87 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases.
+FAVOR-GPT li2024favorgpt utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases.
 Lu et al.
-18 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts.
-RAG-HPO 72 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 81 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information.
-By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 18 and phenotypic and genotypic analysis 72, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
+Lu2025BoostingGPT developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts.
+RAG-HPO garcia2024deepphenotyping connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM Fukushima2025 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information.
+By connecting genomic findings with clinical knowledge through appropriate retrieval mechanisms, these RAG systems support the clinical use of genetic information for tasks such as disease inference Lu2025BoostingGPT as well as phenotypic and genotypic analysis garcia2024deepphenotyping, thereby enabling more personalized treatment decisions in the emerging era of genomic medicine.
 Personalized Health Management
-Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 227.
+Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status Quazi2022HCI.
 This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
 Recent research has produced several personalized health management systems employing different knowledge retrieval strategies.
-HEALIE 29 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients.
-RISE 85 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements.
-A Dual RAG System 50 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations.
-RAP 92 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions.
+HEALIE Kakalou2024 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients.
+RISE wang2024diabetesrag implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements.
+A Dual RAG System Lee2024 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations.
+RAP azimi2025evaluation utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions.
 These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
 Healthcare Education
-Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 228.
+Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences Lemmetty2006HIS.
 These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
-ChatENT 89 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS).
-EyeGPT 153 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries.
+ChatENT long2024chatent leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS).
+EyeGPT chen2024eyegpt transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries.
 Sevgi et al.
-186 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval.
-In nursing education, Zhao et al 229 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification.
-By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.
+sevgi2024medical developed targeted ophthalmology applications, including EyeTeacher for generating educational questions and EyeAssistant for clinical support.
+Both systems are enhanced by integrating clinical guidelines and peer-reviewed documents as retrieval sources.
+In nursing education, Zhao et al ZHAO2025104284 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification.
+By connecting learning objectives with relevant medical knowledge through appropriate retrieval mechanisms, these RAG-based education systems help support healthcare practice and skill training.
 Clinical Research
 Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations.
-230.
+ge2024leveraging.
 Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches.
-PodGPT 34 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development.
-LITURAt 62 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification.
-These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
+PodGPT jia2024podgpt implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development.
+LITURAt Peasley2025 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification.
+These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and supporting the translation of research findings into clinical practice.
+Discussion
+In this section, we summarize the key observations drawn from technological approaches, datasets, evaluation resources, and biomedical applications, examines the limitations of existing work, and further discusses the major challenges faced by RAG systems in the biomedicine domain while outlining promising directions for future research and development.
+Summary of technological approaches
+In Section.III, we reviewed three key technological foundations of biomedical RAG systems identified in our survey: retrieval, reranking, and generation.
+Retrieval methods span traditional sparse models such as BM25bm25, multiple classes of dense retrievers—including general PLMs, commercial embedding APIs, domain-adapted biomedical models, LLM-based embeddings, and specialized architectures such as MedCPTjin2023medcpt and BMRetrieverxu2024bmretriever, and multimodal frameworks.
+Knowledge-graph–augmented retrieval further enhances precision by leveraging structured biomedical ontologies and explicit entity–relation pathways.
+Benefiting from its efficiency and proven generalizability, BM25bm25 remains the leading approach for retrieval; however, it is limited to textual resources.
+Emerging image-based and multimodal systems remain limited, partly because single-vector indexingXia2024RULE, pmlr-v219-ranjit23a, Sun2024FactAware—widely used in text retrieval—may underrepresent the heterogeneous characteristics of multimodal data, making their indexing and retrieval methodologies for multimodal RAG still worthy of further exploration.
+Reranking techniques provide a second-stage selection mechanism, aiming to enhance the quality of the initial rank with more advanced and algorithmically complex strategies.
+Existing systems employ statistical approaches (e.g., RRF, MMR), transformer-basedjia2024podgpt, adejumo2024retrieval and cross-encoder modelsjin2023medcpt, LLM-based relevance classifiersdarnell2024biomedical, craig2024lmrac, as well as domain-specific rerankers designed to better capture biomedical semantics.
+The main challenge for reranking models is generalization at inference time, as they often rely on annotated samples.
+Creating application-specific rank–gold labels for existing datasets is expensive, while automatic annotation carries risks of reduced accuracy, forcing developers to carefully balance this trade-off in practice.
+Lastly, on the generation side, both open-source and commercial LLMs are used extensively, alongside an increasing number of biomedical-specialized models (see Table LABEL:tab:llm_taxonomy), to generate content benefiting from the retrieved biomedical resources.
+We also find that domain adaptation—whether applied to retrievers, rerankers, or generators—enhances terminology handling, conceptual grounding, and factual reliability.
+Building on these observations, commercial LLMs are easy to use and can deliver competitive performance, whereas open-source models—especially biomedical-specialized ones—allow continued training or fine-tuning for specific tasks and even modifications to the model architecture, enabling more flexible application across diverse tasks.
+However, this flexibility also imposes additional computational resource demands on developers.
+For the future development of RAG, we believe that continued progress in multimodal retrieval and knowledge graph integration will expand the scope of information accessible to RAG systems and improve their ability to generate contextually informed, clinically aligned outputs.
+This underscores the growing importance of cross-modal fusion strategies and targeted domain adaptation in advancing biomedical RAG capabilities.
+Summary of dataset and evaluation approaches
+In Sec.IV, we reviewed datasets relevant to the development and evaluation of biomedical RAG systems, categorizing them into Knowledge Source Datasets (Table LABEL:tab:med_knowledge) and Medical Task Datasets (Table LABEL:tab:biomedical-comprehensive).
+Biomedical knowledge sources are highly diverse, and this heterogeneity poses significant challenges for retrieval, as different sources often exhibit distinct linguistic styles and vocabulary patterns.
+On the evaluation side, current benchmarks mainly rely on various Medical Task Datasets, with a particular focus on knowledge-intensive tasks such as Multiple-Choice Examinations (e.g., MedQA jin2020disease, MedMCQA pal2022medmcqa), which simulate medical licensing assessments, and a range of Question Answering (QA) tasksBenAbacha-BMC-2019, singhal2022large, hou2023geneturing, Lozano2024ClinfoAI, ALONSO2024102938.
+However, existing evaluations still lack tasks that are more concrete and aligned with real-world clinical scenarios.
+We also observe a strong dominance of English-language datasets and a heavy reliance on textual benchmarks.
+Substantial future efforts are needed to build richer cross-lingual and multi-modal evaluation resources that better reflect the diversity and complexity of biomedical practice.
+For evaluation metrics, we follow the retrieval–reranking–generation pipeline and review the metrics used at each stage accordingly.
+Retrieval and reranking are typically assessed using standard information-retrieval measures such as recall, precision, nDCG, and MRR, while generation metrics vary widely depending on the specific downstream task.
+To provide a clearer overview, we group generation-level evaluations into four major task types, each associated with its own metric family.
+Although many automatic metrics exist, they often struggle to capture the nuanced reasoning quality required in complex biomedical dialogue and decision-support scenarios metric-survey.
+As a result, human evaluation remains essential.
+However, due to the substantial cost of expert annotation, LLM-based evaluators are increasingly adopted Croxford2025, metric-survey.
+These models can follow instructions and emulate human evaluation criteria, thereby offering a more scalable alternative to manual assessment, although they may also introduce risks such as hallucination and evaluation bias.
+Therefore, developing evaluation frameworks that are both low-cost and highly accurate remains an important direction for future research.
+Summary of RAG applications in biomedicine
+Biomedical RAG applications (Sec.
+IV) span a wide range of healthcare domains—from clinical decision support and precision medicine to medical education and clinical research—highlighting RAG’s capacity to enhance factual grounding, interpretability, and domain specificity for AI-driven healthcare solutions.
+Across the surveyed literature, RAG-based systems have demonstrated utility in various medical specialties, including oncologyLammert2024ExpertLLM, radiologysloan2024automated, ophthalmologychen2024eyegpt, geneticsli2024favorgpt, Lu2025BoostingGPT, Fukushima2025, nursingZHAO2025104284, and rare disease diagnosisZHAO2025104284, Zelin2024RareDisease, garcia2024deepphenotyping.
+These domain-specific implementations suggest that retrieval-augmented approaches can help connect LLMs with biomedical evidence, thereby supporting grounded question answering, enhancing the consistency of clinical reasoning processes, and contributing to clearer and more interpretable outputs for both patient-facing and clinician-facing use cases.
+However, the surveyed applications also reveal several practical constraints.
+Retrieval steps often introduce latency and increase token length, which can reduce system efficiency and limit scalability in real clinical workflowsAmmar2025.
+In addition, current retrievers still struggle with fine-grained biomedical semantics—such as distinguishing closely related clinical entities or integrating cross-modal evidence—which can lead to suboptimal grounding even when high-quality sources existTan2024.
+Addressing these issues will be important for deploying RAG systems in settings that demand both timely responses and high diagnostic precision.
 Challenges and Future Direction
 Trustworthiness
-Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 1, 231.
+Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans Fan2024RAG, deng2024pandora.
 Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
 Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns.
-Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 181.
-Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications.
+Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source gao2023retrieval.
+Although certain re-ranking methodologies described in Section 3.2 can partially mitigate these issues, significant challenges remain in biomedical RAG applications.
 Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents.
 A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources.
-For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 139.
+For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications lee2024enhancing.
 This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems.
 Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process.
-Due to the uneven distribution of demographic attributes in training corpora 2, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases.
+Due to the uneven distribution of demographic attributes in training corpora he2023survey, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases.
 Addressing the fairness concerns also represents an important direction for future research.
@@
 This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets.
-The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 232, 233.
+The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored neveol2018clinical, Englishllm.
 This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
@@
 Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms.
-Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 234, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
+Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages mondshine2025beyond, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
 Multimodality
 Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g.
-X-ray 203, 235, 236), time series (e.g.
-ECG 152), sound recordings (e.g.
-lung sounds recorded 237), and video 238.
+X-ray johnson2019mimic, pelka2018roco, lin2023pmc), time series (e.g.
+ECG pmlr-v225-yu23b), sound recordings (e.g.
+lung sounds recorded fraiwan2021dataset), and video gupta2023dataset.
 Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored.
 Multi-modal data presents several intractable challenges for retrieval systems.
-For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 239.
+For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes zhong2024class.
 Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression.
@@
 The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios.
-Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 1.
+Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge Fan2024RAG.
 However, these models typically require high-performance computing clusters to operate effectively.
-For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 240.
-Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 241, 242.
+For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs deepseekv3technicalreport.
+Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks brown2020language, hoffmann2022training.
 Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive.
 Several promising approaches have emerged to address these constraints.
-One solution involves developing smaller, distilled LLMs derived from their larger counterparts 243.
-Models such as QWQ 160 and Mistral Small 3.1 40 demonstrate competitive performance despite having significantly reduced parameter counts.
-Applying Portable LLMs 244 as generation models is also a promising method.
+One solution involves developing smaller, distilled LLMs derived from their larger counterparts xu2024survey.
+Models such as QWQ qwen32b_2024 and Mistral Small 3.1 mistral_small_24b_2024 demonstrate competitive performance despite having significantly reduced parameter counts.
+Applying portable LLMs khan2024portllm as generation models is also a promising method.
 Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services.
@@
 Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes.
-During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 245.
-Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 246 and the Health Insurance Portability and Accountability Act 247.
+During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential meystre2010automatic.
+Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation GDPR2016 and the Health Insurance Portability and Accountability Act HIPAA1996.
 Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development.